### v1

# Mutual Information Estimation via Score-to-Fisher Bridge for Nonlinear Gaussian Noise Channels

Abstract
We present a numerical method to evaluate mutual information (MI) in nonlinear Gaussian noise channels by using denoisingscore matching learning for estimating the score function of channel output. Via de Bruijn’s identity and the I–MMSE relation, Fisher information estimated from the learned score yields accurate estimates of MI through an integral representation of MI for a variety of priors and channel nonlinearities (e.g., elementwise nonlinearity). In this work, we propose a comprehensive theoretical foundation for the Score-to-Fisher bridge methodology, along with practical guidelines for its implementation. We also conduct extensive validation experiments, comparing our approach with closed-form solutions and a kernel density estimation baseline. The results of our numerical experiments demonstrate that the proposed method is both practical and efficient for mutual information estimation in nonlinear Gaussian noise channels.
I Introduction
We consider additive Gaussian noise channels with a deterministic nonlinear front-end,
$Y_{t}=f(X)+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (1)
and study how to estimate the mutual information $I(X;Y_{t})$ efficiently and accurately. The nonlinear front-end can be represented by a deterministic function $f$. Such models arise widely due to nonlinear front-end saturation/clipping, DAC/ADC nonlinearity, photonic modulators, optical fibers, or learned encoders implemented with neural networks. The value of the mutual information $I(X;Y_{t})$ is important for evaluating the ultimate performance of coded communication systems and it provides useful information for designing the system such as coding, preprocessing, or decoding/detection algorithms.
When $f$ is nonlinear, evaluating $I(X;Y_{t})$ is, in general, challenging in practice because accurate posterior means $\mathbb{E}[X|Y_{t}]$ and normalized likelihoods are inaccessible or computationally intractable. For linear Gaussian channels, the de Bruijn identity and the I–MMSE relation yield
$\frac{d}{dt}I(X;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}=-\frac{1}{2t^{2}}\mathrm{mmse}(t),$ (2)
so that the mutual information $I(X;Y_{t})$ can be obtained by integrating either the Fisher information $J(Y_{t})$ or the MMSE. Our key observation is that the deterministic nonlinearity can be absorbed into a transformed input $W=f(X)$, so that
$I(X;Y_{t})=I(W;Y_{t}),\quad Y_{t}=W+Z_{t},$
holds for any deterministic function $f$. Hence, for general $f$ we can still rely on AWGN identities:
$\mathrm{mmse}_{W}(t)=nt-t^{2}J(Y_{t}),$
and consequently obtain $I(X;Y_{t})$ by integrating $J(Y_{t})$ or $\mathrm{mmse}_{W}(t)$.
In this paper, we propose a score-based estimator of mutual information tailored to $Y_{t}=f(X)+Z_{t}$. The method learns the score of channel output, or the marginal $p_{Y_{t}}$,
$s_{Y_{t}}(\bm{y})\equiv\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y}),$ (3)
using denoising score matching (DSM) and then obtain an estimate of the Fisher information $\widehat{J}(Y_{t}).$ From this, we obtain an estimate of the mutual information via an integral representation of MI:
$\widehat{I}(X;Y_{T})=\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-\widehat{J}(Y_{t})\Big)\,dt.$ (4)
Crucially, the approach never requires access to the posterior $p_{X|Y}(\bm{x}|\bm{y})$. All we need are forward samples $\bm{x}\sim p_{X}$, evaluations of $\bm{w}=f(\bm{x})$, and Gaussian noise addition, which enables a single-loop estimator of MMSE and the Fisher information.
The main contributions of this paper are summarized as follows:
  • Score-only MI estimation for nonlinear front-ends. We formalize a pipeline— denoising score matching of $p_{Y_{t}}$ $\Rightarrow$ Fisher estimation $\Rightarrow$ MMSE/MI integration—that applies to arbitrary deterministic $f$ as soon as $p_{Y_{t}}$ can be sampled.
  • Unified theoretical grounding. Leveraging de Bruijn and I–MMSE, we derive $\mathrm{mmse}_{W}(t)=nt-t^{2}J(Y_{t})$ and connect $I(X;Y_{t})$ to $I(W;Y_{t})$ with $W=f(X)$, extending linear-Gaussian identities to general nonlinear front-ends without ambiguity.
  • Practical recipe. We detail implementation choices that make the estimator stable and reproducible: logarithmic grids in $t$, tail corrections for the MI integral, and a diagnostic “double estimate” that compares a Tweedie-based direct MMSE with the Fisher-based MMSE.
  • Validation and use-cases. We verify the estimator on Gaussian inputs where closed forms are available, and report experiments on composite nonlinear channels.
In linear Gaussian channels, I–MMSE and multivariate de Bruijn provide differential and integral representations of information quantities have been extensively studied. For nonlinear $f$, however, posterior-free and normalization-free MI estimators remain limited. Building on score matching and denoising-based score learning, our approach attaches directly to the marginal $p_{Y_{t}}$, circumventing posterior computations and partition functions. The estimate of mutual information can be used many applications such as design of coding systems, optimization of preprocessing, and design of detection and decoding algorithms.
II Preliminaries
II-A Notation
Unless otherwise stated, all logarithms are natural (nats). Bold letters denote vectors in $\mathbb{R}^{n}$. For a random vector $U$ with density $p_{U}$, we write $h(U)$ for its differential entropy, $I(\cdot;\cdot)$ for mutual information, and $\operatorname{Cov}(U)$ for its covariance matrix. The (scalar) Fisher information is defined via the score function $s_{U}(\bm{u})\equiv\nabla_{\bm{u}}\log p_{U}(\bm{u})$ as
$J(U)\equiv\mathbb{E}\!\left[\|\nabla_{\bm{u}}\log p_{U}(\bm{u})\|^{2}\right]\quad\text{with }\bm{u}\sim p_{U}$ (5)
see, e.g., [1, 2].
II-B Related works
For channels with memory, simulation-based methods have been developed to evaluate information rates without closed-form expressions. Arnold et al. [3] proposed a simulation-based framework that computes upper and lower bounds on the information rate by combining auxiliary finite-state channel models with forward–backward recursions and importance sampling; the method is broadly applicable to ISI and finite-state channels where exact evaluation is infeasible. Building on this line, Sadeghi et al. [4] optimized these bounds over the choice of auxiliary models, yielding systematically improved estimates and a practical tool for design and benchmarking of channels with memory.
Another line of work estimates $I(X;Y)$ directly from its definition via simulation-based plug-in estimators: one approximates $p_{Y}$ (or $h(Y)$) using Kernel Density Estimation (KDE) or mixtures [5] or employs $k$-nearest neighbors (kkNN) estimators [6]. In the additive Gaussian setting, recent theory shows that plug-in estimation of $h(X+Z)$ is in fact optimal [7].
II-C AWGN channel and score function
Let $X\in\mathbb{R}^{n}$ be a channel input random vector with known prior probability density function (PDF) $p_{X}$, and consider the AWGN channel family:
$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),\quad t>0,$ (6)
where $t$ represents the noise variance. Let $p_{Y_{t}}$ denote the marginal density of $Y_{t}$. The score function of $p_{Y_{t}}$ is defined by $s_{Y_{t}}(\bm{y})\equiv\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y}).$ Throughout, we assume that sampling from $p_{X}$ is computationally feasible by forward sampling or Langevin sampling using the score $\nabla_{\bm{x}}\log p_{X}(\bm{x})$ (e.g., Rossky et al. [8], Welling and Teh [9]).
The Fisher information of $Y_{t}$ is given by
$J(Y_{t})\equiv\mathbb{E}_{Y_{t}}\bigl[\|s_{Y_{t}}(\bm{y})\|^{2}\bigr].$ (7)
Under standard regularity conditions, the following Tweedie identity [10, 11, 12] holds for $t>0$:
$\mathbb{E}[X|Y_{t}=\bm{y}]=\bm{y}+ts_{Y_{t}}(\bm{y}).$ (8)
II-D I–MMSE relation
For the SNR-parameterized AWGN model
$Y_{\gamma}=\sqrt{\gamma}X+N,\quad N\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n}),\quad N\perp X,$ (9)
the I–MMSE relation [13, 14] states that the identity
$\frac{d}{d\gamma}I(X;Y_{\gamma})=\frac{1}{2}\mathrm{mmse}(\gamma),$ (10)
holds for any snr $\gamma>0$ where $\mathrm{mmse}(\gamma)$ is defined by
$\mathrm{mmse}(\gamma)\equiv\mathbb{E}\bigl[\|X-\mathbb{E}[X|Y_{\gamma}]\|^{2}\bigr].$ (11)
With the change of variables $\gamma=1/t$ (and noting $I(X;Y_{\gamma})=I(X;Y_{t})$), an equivalent form for our variance parameterization $Y_{t}=X+Z_{t}$ is given by
$\frac{d}{dt}\,I(X;Y_{t})=-\frac{1}{2t^{2}}\mathrm{mmse}(t),$ (12)
where $\mathrm{mmse}(t)\equiv\mathbb{E}\bigl[\|X-\mathbb{E}[X|Y_{t}]\|^{2}\bigr].$
II-E de Bruijn identity
Let $h(Y_{t})$ denote the differential entropy of $Y_{t}$. Then the (multivariate) de Bruijn identity gives
$\frac{d}{dt}h(Y_{t})=\frac{1}{2}J(Y_{t}).$ (13)
The de Bruijn identity was discussed in [15, 16].
II-F Denoising score matching (DSM) method
DSM [17, 18, 19, 20] provides a practical way to learn the score function of a Gaussian-smoothed distribution without computing any normalization constants or partition functions. Let $p$ be a target density on $\mathbb{R}^{n}$ and define a Gaussian-smoothed sample:
$\bm{v}=\bm{w}+\sigma\bm{\epsilon},\quad\bm{w}\sim p,\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n}).$ (14)
Writing $p_{\sigma}\equiv p*\phi_{\sigma^{2}}$ for the convolution with a zero-mean Gaussian PDF $\phi_{\sigma^{2}}$ with variance $\sigma^{2}$, the DSM loss function for a parametric score model $s_{\bm{\theta}}:\mathbb{R}^{n}\!\to\!\mathbb{R}^{n}$ ($\bm{\theta}$ represents the vector of adjustable parameters) is given by
$\mathbb{E}_{\bm{w},\bm{\epsilon}}\Big[\big\|s_{\theta}(\bm{v})+\epsilon/\sigma\,\big\|^{2}\Big]$ (15)
$=\mathbb{E}_{\bm{v}\sim p_{\sigma}}\Big[\big\|\,s_{\theta}(\bm{v})-\nabla_{\bm{v}}\log p_{\sigma}(\bm{v})\,\big\|^{2}\Big]+\text{const},$ (16)
where this equivalence was shown by Vincent [18]. This means that minimizing the DSM loss function is equivalent to minimizing the expectation of the squared distance between the learned score function and the true score function, which is sometimes called the Fisher divergence. This means that the unique minimizer of the DSM loss function satisfies $s_{\bm{\theta}}(\bm{y})\simeq\nabla_{\bm{y}}\log p_{\sigma}(\bm{y})$ in high probability. Equivalently, since $\bm{v}-\bm{w}=\sigma\bm{\epsilon}$, one may implement the loss function as $(\bm{v}-\bm{w})/\sigma^{2}$:
$\mathbb{E}_{\bm{w},\bm{\epsilon}}\big[\|s_{\theta}(\bm{v})+(\bm{v}-\bm{w})/\sigma^{2}\,\|^{2}\big],$ (17)
which is an equivalent form of the DSM loss function.
In practice, we use a neural network to parameterize the score model $s_{\bm{\theta}}(\bm{v})$. Such a score model is also called a score network. A mini-batch training with stochastic gradient descent or other variants such as Adam is typically used to minimize the DSM loss function.
III Estimating Mutual Information in AWGN Channels
In this section, we present a method to estimate the MMSE and MI for AWGN channels described by
$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),\quad t>0,$ (18)
where $X\sim p_{X}$. The results obtained in this section will be extended to general nonlinear channels in Section V.
III-A MMSE–Fisher relation
For any Gaussian channel, the following proposition holds.
Proposition 1.
For all $t>0$, the identity
$\mathrm{mmse}(t)=nt-t^{2}J(Y_{t}).$ (19)
holds [13, 14, 21].
Proof.
A sketch of proof is the following. Using $I(X;Y_{t})=h(Y_{t})-h(Z_{t})$ with $h(Z_{t})=(n/2)\log(2\pi e\,t)$, we have
$\frac{d}{dt}I(X;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}$ (20)
by de Bruijn identity. Equating this with the I–MMSE form
$\frac{d}{dt}I(X;Y_{t})=-\frac{1}{2t^{2}}\mathrm{mmse}(t)$ (21)
gives $\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$. ∎
Reeves et al. [21] investigated this identity in the context of linear Gaussian channels, highlighting its implications for non-Gaussian input distributions. Their work extends the I–MMSE framework beyond the classical Gaussian setting.
III-B Integral representations of mutual information
From the identity (20) and the fact that $\lim_{t\to\infty}I(X;Y_{t})=0$ (since $Y_{t}$ and $X$ become independent in the limit), integrating over $t\in[T,\infty)$ yields
$I(X;Y_{T})=$ $\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-J(Y_{t})\Big)\,dt$ (22)
$=$ $\frac{1}{2}\int_{T}^{\infty}\!\frac{\mathrm{mmse}(t)}{t^{2}}\,dt.$ (23)
The integrand decays as $O(1/t^{2})$, since
$0\ \leq\ \frac{n}{t}-J(Y_{t})\;=\;\frac{\mathrm{mmse}(t)}{t^{2}}\ \leq\ \frac{\operatorname{tr}\operatorname{Cov}(X)}{t^{2}}.$ (24)
The integral representation directly follows by combining the de Bruijn identity [15] with the I–MMSE relation [13, 14].
III-C Estimation of Fisher Information
Using DSM method, we can estimate the score function of $Y_{t}$. With $p=p_{X}$ and $\sigma^{2}=t$, the learned score model coincides with the score of $Y_{t}=X+Z_{t}$:
$s_{\bm{\theta}}(\cdot)\approx\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y})=s_{Y_{t}}(\bm{y}),$ (25)
where $s_{\bm{\theta}}(\cdot)$ can be implemented with a neural network with parameter vector $\bm{\theta}$. This enables estimation of $J(Y_{t})=\mathbb{E}\|s_{Y_{t}}(Y_{t})\|^{2}$ from samples of $\bm{x}\sim p_{X}$ and Gaussian noises.
In particular, for a fixed noise level $t$, the DSM loss function can be written as
$\mathcal{L}_{t}(\bm{\theta})\;\equiv\;\mathbb{E}_{\bm{x},\bm{\epsilon}}\!\left[\left\|\,s_{\bm{\theta}}(\bm{x}+\sqrt{t}\,\bm{\epsilon})+\bm{\epsilon}/\sqrt{t}\,\right\|^{2}\right]$ (26)
$=\mathbb{E}_{Y_{t}}\!\left[\left\|\,s_{\bm{\theta}}(Y_{t})-s_{Y_{t}}(Y_{t})\,\right\|^{2}\right]+\text{const},$ (27)
where the constant term does not depend on $\bm{\theta}$. Hence, with a sufficiently rich function class for $s_{\bm{\theta}}$ (e.g., a universal neural network) and sufficiently many samples, any sequence of empirical minimizers achieves $\mathbb{E}_{Y_{t}}\!\left[\left\|\,s_{\bm{\theta}}(Y_{t})-s_{Y_{t}}(Y_{t})\,\right\|^{2}\right]\rightarrow 0,$ i.e., the learned score converges to the true score in the Fisher divergence sense.
Moreover, with i.i.d. draws $\bm{y}_{i}=\bm{x}_{i}+\sqrt{t}\,\bm{\epsilon}_{i}$ from $p_{Y_{t}}(i=1,2,...,N)$, the Monte Carlo estimator:
$\widehat{J}_{t}\triangleq\frac{1}{N}\sum_{i=1}^{N}\|s_{Y_{t}}(\bm{y}_{i})\|^{2}$ (28)
satisfies $\widehat{J}_{t}\xrightarrow{\text{a.s.}}J(Y_{t})$ as $N\to\infty$ (law of large numbers), and $\sqrt{N}\,(\widehat{J}_{t}-J(Y_{t}))\Rightarrow\mathcal{N}\big(0,\operatorname{Var}(\|s_{Y_{t}}(\bm{y})\|^{2})\big)$ (central limit theorem), so the estimation error decays at rate $O(N^{-1/2})$. In practice we plug in the learned score $s_{\bm{\theta}}$ in place of $s_{Y_{t}}$; under Fisher divergence consistency of $s_{\bm{\theta}}$, the plug-in estimator $(1/N)\sum_{i=1}^{N}\|s_{\theta}(\bm{y}_{i})\|^{2}$ is expected to be also consistent for $J(Y_{t})$.
III-D Gaussian input case for validation
In this subsection we provide closed-form expressions for the Gaussian-input case $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ in the AWGN channel family $Y_{t}=X+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$. It is well known [2] that $I(X;Y_{t})=(n/2)\log(1+P/t)$; see also [13] for the relation to MMSE.
Proposition 2 (Closed forms for Gaussian inputs).
Let $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ and $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$, independent. Then, for all $t>0$, we have the following closed forms:
$\frac{1}{n}I(X;Y_{t})=\frac{1}{2}\log\!\Big(1+\frac{P}{t}\Big),$ (29)
$\frac{1}{n}h(X\mid Y_{t})=\frac{1}{2}\log\!\Big(2\pi e\,\frac{Pt}{P+t}\Big),$ (30)
$\mathrm{mmse}(t)=n\,\frac{Pt}{P+t},\quad J(Y_{t})=\frac{n}{P+t}.$ (31)
Sketch of proof: .
From the Gaussian model $Y_{t}=X+Z_{t}$, it is known that $I(X;Y_{t})=(n/2)\log(1+P/t)$ holds. Moreover $Y_{t}\sim\mathcal{N}(\mathbf{0},(P+t)\mathbf{I}_{n})$, hence we have
$s_{Y_{t}}(\bm{y})=\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y})=-(P+t)^{-1}\bm{y}$ (32)
and $J(Y_{t})=\mathbb{E}\|s_{Y_{t}}(Y_{t})\|^{2}=n/(P+t)$. The remaining identities follow by substituting $J(Y_{t})$ into $\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$ and by $h(X\mid Y_{t})=h(Y_{t})-I(X;Y_{t})$ with $h(Y_{t})=(n/2)\log\big((2\pi e)(P+t)\big).$
III-E Linear Gaussian channel case
Consider a linear Gaussian channel model:
$Y_{t}=AX+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (33)
where $A\in\mathbb{R}^{n\times n}$ is a square, invertible matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$. This setting preserves the Gaussian structure while introducing nontrivial correlation through the linear mixing due to $A$.
Proposition 3 (Closed forms for linear Gaussian channels).
Let $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$, $A\in\mathbb{R}^{n\times n}$ nonsingular, and $Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n})$, independent. Then for all $t>0$:
$I(X;Y_{t})$ $=\tfrac{1}{2}\log\det\!\Bigl(I_{n}+\tfrac{P}{t}AA^{\top}\Bigr),$ (34)
$\mathrm{mmse}(t)$ $=\mathrm{tr}\!\left(\Bigl(P^{-1}I_{n}+\tfrac{1}{t}A^{\top}A\Bigr)^{-1}\right),$ (35)
$J(Y_{t})$ $=\mathrm{tr}\!\Bigl((PAA^{\top}+tI_{n})^{-1}\Bigr).$ (36)
Sketch of proof:.
Since $AX$ is Gaussian with covariance $PAA^{\top}$, we have $Y_{t}\sim\mathcal{N}(\mathbf{0},PAA^{\top}+tI_{n})$. The mutual information (34) follows from the determinant formula for Gaussian differential entropy [2, 14]. The Fisher information (36) is obtained directly from the Gaussian score function $s_{Y_{t}}(\bm{y})=-(PAA^{\top}+tI_{n})^{-1}\bm{y}$. The MMSE formula (35) comes from the linear MMSE estimator of $X$ given $Y_{t}$. ∎
IV Estimating MMSE and Mutual Information
Using the identity between $\mathrm{mmse}(t)$ and the Fisher information, we can evaluate $I(X;Y_{T})$ by numerical integration. In this section, we present details of the method to estimate the MMSE and MI for the AWGN channel family.
IV-A Score-based Estimation Method for MI and MMSE
The proposed estimation algorithm is summarized in Algorithm 1. The core of the algorithm is the DSM learning for $p_{Y_{t}}$ and the estimation of the Fisher information from the score model.
Algorithm 1 Score-to-Fisher bridge: Estimation Algorithm for MI and MMSE
1: Prepare a logarithmically spaced grid over $t$ (details of the log grid are described in Subsection IV-C).
2: Sample generation: Generate a random mini-batch $\mathcal{B}$ consisting of multiple pairs of $(\bm{x},\bm{\epsilon})$ for mini-batch training with forward sampling of $X$: $\bm{x}\sim p_{X},\ \bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{n}).$
3: Denoising score matching (DSM) learning for $p_{Y_{t}}$: train the score model parameters by minimizing the empirical DSM loss:
$\sum_{(\bm{x},\bm{\epsilon})\in\mathcal{B}}\big\|\,s_{\bm{\theta}}(\bm{x}+\sqrt{t}\,\bm{\epsilon})\;+\;\bm{\epsilon}/\sqrt{t}\,\big\|^{2}.$ (37)
Any stochastic optimizer, such as Adam, can be applied to minimized the empirical loss function.
4: Estimating the Fisher information: for each $t$, generate $\bm{y}_{1},\ldots,\bm{y}_{N}$ (e.g., by forward sampling $\bm{y}_{i}=\bm{x}_{i}+\sqrt{t}\,\bm{\epsilon}_{i}$), and compute the estimated Fisher information:
$\widehat{J}_{t}\equiv\frac{1}{N}\sum_{i=1}^{N}\big\|s_{\theta}(\bm{y}_{i})\big\|^{2},$ (38)
as an approximation of $J(Y_{t})$.
5: Estimating the MMSE:
$\widehat{\mathrm{mmse}}(t)=nt-t^{2}\widehat{J}_{t}.$ (39)
6: Estimating the MI via the integral representation:
$\widehat{I(X;Y_{T})}=\frac{1}{2}\int_{T}^{\infty}\Big(\frac{n}{t}-\widehat{J}_{t}\Big)\,dt,$ (40)
evaluated numerically.
In the above algorithm, we consider two alternatives for the DSM learning:
  • Scheme A (per-$t$ training). For each fixed $t$, minimize
    $\mathbb{E}\big\|s_{\theta}(\bm{x}+\sqrt{t}\bm{\epsilon})+\bm{\epsilon}/\sqrt{t}\big\|^{2}$
    to obtain the learned score model.
  • Scheme B (noise-conditional model). Learn $s_{\theta}(\bm{y},t)$ by minimizing
    $\mathbb{E}_{t\sim\mathcal{T}}\,\lambda(t)\ \mathbb{E}_{\bm{x},\bm{\epsilon}}\,\big\|s_{\bm{\theta}}(\bm{x}+\sqrt{t}\,\bm{\epsilon},t)+\bm{\epsilon}/\sqrt{t}\big\|^{2},$ (41)
    where $\mathcal{T}$ is a training distribution over $t$ (e.g., logarithmically spaced as described later) and $\lambda(t)$ is a weight (e.g., $\lambda(t)=t$).
In the case of Scheme B, the score network has two inputs: the channel output $\bm{y}$ and the noise level $t$. In generative modeling frameworks for image generation, Scheme B is often used to train a noise-conditional score model. The training distribution $\mathcal{T}$ is logarithmically spaced as described in Subsection IV-C. The weight $\lambda(t)$ is a weight function that is used to weight the loss function.
IV-B Validating the Fisher-based estimate
Although Algorithm 1 presented in the previous subsection is an asymptotically consistent method to estimate the MMSE and MI, the accuracy of the estimate is not clear in the case of finite samples. One possible source of the error is the estimation error of the Fisher information due to inaccurate estimation of the score model. Although we can expect the accurate score matching in the Fisher divergence sense in the case of infinite samples, the convergence rate depends on the prior distribution of the source. We also have to consider the numerical error of the numerical integration.
In this section, we present a method to validate the estimated MMSE and MI. As a diagnostic purpose, we can compute the direct MMSE estimate based on the Tweedie formula:
$\widehat{\mathrm{mmse}}_{\rm dir}(t)\equiv\frac{1}{N}\sum_{i=1}^{N}\big\|\bm{x}_{i}-\{\bm{y}_{i}(t)+t\,s_{\theta}(\bm{y}_{i}(t))\}\big\|^{2},$ (42)
and check agreement with the Fisher-based estimate:
$\widehat{\mathrm{mmse}}_{F}(t)\equiv nt-t^{2}\widehat{J}_{t}.$ (43)
This consistency check is useful to validate the accuracy of the Fisher-based estimate.
IV-C Log-domain trapezoid integration
The integrand in the integration representation of MI is smooth and may have a long tail. The choice of the integration grid is critical for achieving high numerical accuracy. In this subsection, we present a geometric (log-spaced) grid for the integration. For $t\in[t_{\min},t_{\max}]$ with $M$ points, take $\log t$ to be equally spaced:
$t_{k}=t_{\min}\left(\frac{t_{\max}}{t_{\min}}\right)^{\frac{k}{M-1}},\qquad k=0,1,\ldots,M-1.$ (44)
Equivalently in standard deviation, $\sigma_{k}\equiv\sqrt{t_{k}}$ with $\sigma_{k}=\sigma_{\min}(\sigma_{\max}/\sigma_{\min})^{k/(M-1)}$.
Rather than applying the trapezoid rule directly on the non-uniform $t$-grid, we perform the change of variables $u=\log t$ to obtain uniform spacing in the log domain. Using the integral representation
$I(X;Y_{T})=\frac{1}{2}\int_{T}^{\infty}\!\left(\frac{n}{t}-J(Y_{t})\right)\,dt,$ (45)
and substituting $u=\log t$ (so $dt=e^{u}du$ and $t=e^{u}$), we obtain
$I(X;Y_{T})=\frac{1}{2}\int_{\log T}^{\infty}\!\Big(n-e^{u}\,J(Y_{e^{u}})\Big)\,du.$ (46)
Since the $t$-grid (44) corresponds to uniform spacing in $u=\log t$ with step size $\Delta_{u}\equiv\log(t_{\max}/t_{\min})/(M-1)$, we apply the standard trapezoid rule (uniform spacing):
$\widehat{I}(X;Y_{T})\approx\frac{1}{2}\sum_{k=k_{T}}^{M-2}\frac{\ell(u_{k})+\ell(u_{k+1})}{2}\,\Delta_{u}\;+\;\text{tail},$ (47)
where $\ell(u)\equiv n-e^{u}\,\widehat{J}(Y_{e^{u}})$ (equivalently $\ell(t)=n-t\,\widehat{J}(Y_{t})$), and $k_{T}$ is the smallest index with $t_{k_{T}}\geq T$. This log-domain integration offers several benefits over direct $t$-domain trapezoid integration: (1) uniform grid spacing reduces discretization error, (2) the transformation naturally handles a wide dynamic range in $t$, and (3) convexity bias is significantly reduced. In our experiments, this approach improved MI estimation accuracy by over an order of magnitude.
The tail contribution above $t_{\max}$ can be estimated using the asymptotic behavior. For Gaussian inputs, $\mathrm{mmse}(t)\to\operatorname{tr}\mathrm{Cov}(X)$ as $t\to\infty$, giving
$\text{tail}\approx\frac{1}{2}\int_{t_{\max}}^{\infty}\frac{\mathrm{mmse}(t)}{t^{2}}\,dt\approx\frac{1}{2}\,\frac{\operatorname{tr}\mathrm{Cov}(X)}{t_{\max}}.$ (48)
The trace $\operatorname{tr}\mathrm{Cov}(X)$ can be estimated from samples of $p_{X}$ and used as a tail correction.
While the tail correction formula (48) is exact for Gaussian inputs, its leading-order term $\mathrm{tr}\,\mathrm{Cov}(X)/t_{\max}$ remains valid for arbitrary inputs, since $\mathrm{mmse}(t)\to\mathrm{tr}\,\mathrm{Cov}(X)$ as $t\to\infty$. Therefore, the same tail correction can be applied as a universal approximation, with the understanding that higher-order terms may depend on the non-Gaussianity of $X$.
V Nonlinear channel
Let us consider our primal target channel:
$Y_{t}=f(X)+Z_{t},$
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a deterministic function. The mutual information equivalence presented in the following proposition is a key observation which enables the score-based estimation to be applied to general nonlinear channels.
V-A Mutual information equivalence
Let $W\equiv f(X)$ be a measurable deterministic mapping, and consider $Y_{t}=W+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n})$ independent of $(X,W)$. Then the conditional law of $Y_{t}$ given $X$ depends on $X$ only through $W$. This means that $X\to W\to Y_{t}$ forms a Markov chain and $Y_{t}\perp X\mid W$.
We can prove mutual information equivalence in the following proposition.
Proposition 4.
Let $Y_{t}=W+Z_{t}(t>0)$ with $Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n})$ independent of $(X,W)$ and $W=f(X)$, where $f$ is a measurable deterministic function. Then we have
$I(X;Y_{t})=I(W;Y_{t}).$ (49)
Proof.
By the chain rule of mutual information, we have
$I(X;Y_{t})=I(W;Y_{t})+I(X;Y_{t}\mid W)-I(W;Y_{t}\mid X).$ (50)
Here $I(X;Y_{t}\mid W)=0$ because of the conditional independence $Y_{t}\perp X\mid W$, and $I(W;Y_{t}\mid X)=0$ holds because $W$ is a function of $X$. Hence $I(X;Y_{t})=I(W;Y_{t})$. An equivalent proof uses the data processing inequality twice: $X\to W\to Y_{t}$ gives $I(X;Y_{t})\leq I(W;Y_{t})$, while $W\to X\to Y_{t}$ (since $W=f(X)$) gives $I(W;Y_{t})\leq I(X;Y_{t})$, yielding equality. ∎
From Proposition 4, we have the MI equivalence $I(X;Y_{t})=I(W;Y_{t})$, which implies that the score-based estimation method presented in the previous section can be applied to evaluate $I(X;Y_{T})$. Namely, instead of $X$, from the samples of $W$, we can estimate the Fisher information of $Y_{t}$ and the estimate of the Fisher information can be used to evaluate $I(W;Y_{t})$. The equivalence relation $I(X;Y_{t})=I(W;Y_{t})$ enables to obtain the estimate of $I(X;Y_{t})$ from the estimate of $I(W;Y_{t})$. Note that no invertibility of $f$ is required in this estimation method.
V-B Score-based estimation method for nonlinear channels
In the following discussin, we assume that the evalution of $f$ is computationally feasible. Namely, sampling from $p_{W}$ is feasible. Some parts of Algorithm 1 are replaced by the following steps:
  • Sample $W$ by $W=f(X)$, where $X\sim p_{X}$.
  • Estimate the Fisher information of $Y_{t}$ from the samples of $W$ by the score-based estimation method in Algorithm 1.
  • Use the estimated Fisher information to evaluate the integral form of MI.
  • Use the relationship $I(X;Y_{t})=I(W;Y_{t})$ to obtain the estimate of $I(X;Y_{t})$.
The estimation procedure is still a single-loop procedure, which does not require the access to the posterior $p_{X|Y_{t}}$.
VI Numerical experiments
VI-A Gaussian input case for validation
To validate the proposed score-based estimation method, we first consider the Gaussian input case where exact closed-form expressions are available. Specifically, we consider the Gaussian input case $X\sim\mathcal{N}(0,PI_{n})$ with the AWGN channel model $Y_{t}=X+\sqrt{t}Z$, where $Z\sim\mathcal{N}(0,I_{n})$.
VI-A1 Experimental Setup
We implemented the DSM method with the following configuration: The score network is implemented as a fully connected neural network consisting of 3 hidden layers, each containing 128 units with SiLU (Swish) activation functions. The input and output dimensions match the problem dimension $n$, where we test with $n\in\{4,8,16\}$ to evaluate scalability. For the training configuration, we set the signal power to $P=1.0$ and construct a logarithmically spaced grid of $M=10$ noise variance values ranging from $t\in[P/200,200P]$, which covers four orders of magnitude. At each noise level $t$, we train a separate score network for 300 iterations using mini-batches of 4096 samples. The optimization is performed using the Adam optimizer with a learning rate of $10^{-3}$, and gradient clipping is applied with a maximum norm of 1.0 to ensure stable training.
VI-A2 Validating the Fisher Information Estimate
For Fisher information estimation, we generate 100,000 Monte Carlo samples at each noise variance $t$ to ensure statistical reliability. The ground truth values are computed using the analytical formula $J(Y_{t})=n/(P+t)$, which is available for Gaussian input distributions.
Figure 1 presents the estimated Fisher information $\hat{J}(Y_{t})$ compared with the analytical ground truth $J(Y_{t})=n/(P+t)$ for dimensions $n\in\{4,8,16\}$. The results demonstrate excellent agreement between the DSM-based estimates and theoretical values across multiple orders of magnitude in the noise variance $t$.
Figure 1: Fisher information $J(Y_{t})$ versus noise variance $t$ for Gaussian input with $P=1$ and dimensions $n\in\{4,8,16\}$. Solid lines with circular markers show ground truth values $J(Y_{t})=n/(P+t)$, while dashed lines with triangular markers show DSM estimates $\hat{J}(Y_{t})$.
The results confirm that:
  1. 1.
    The DSM-based score estimation accurately captures the Fisher information across a wide range of noise levels ($t\in[0.005,200]$).
  2. 2.
    The method scales well with dimension, maintaining low relative errors even for $n=16$.
  3. 3.
    The median relative error remains below 1.3% for all tested dimensions, validating the theoretical framework.
VI-A3 Validating the Fisher-based MMSE Estimate
To validate our Fisher-based MMSE estimation, we compare three methods using the same experimental setup with $n=4$: (1) true MMSE (analytical ground truth), (2) Tweedie-based direct estimation using
$\widehat{\text{mmse}}_{\text{dir}}(t)\equiv\frac{1}{N}\sum_{i=1}^{N}\|\bm{x}_{i}-(\bm{y}_{i}+ts_{\bm{\theta}}(\bm{y}_{i}))\|^{2},$ (51)
and (3) our Fisher-based method using
$\widehat{\text{mmse}}_{\text{F}}(t)=nt-t^{2}\hat{J}(Y_{t}).$ (52)
The noise variance range is set to $t\in[P/200,50P]$ to avoid numerical instabilities at extremely high noise levels. The results are shown in Figure 2. The two estimates are in excellent agreement with the true MMSE in the wide range of noise variance.
Figure 2: Comparison of MMSE estimation methods for Gaussian input ($n=4$, $P=1$).
VI-A4 Mutual Information Estimation
We demonstrate the complete pipeline from the Score-to-Fisher bridge for mutual information estimation using the integral representation of MI. The experiment uses the same configuration as previous tests with $n=4$ and $P=1.0$, implementing log-domain trapezoid integration with tail correction. Figure 3 presents the estimated mutual information (per symbol) $\hat{I}(X;Y_{t})/n$ compared with the analytical ground truth $I(X;Y_{t})/n=(1/2)\log(1+P/t)$ for Gaussian input. The DSM-based estimates demonstrate excellent agreement with the theoretical values across the entire noise variance range.
Figure 3: Mutual information (per symbol) $I(X;Y_{t})/n$ versus noise variance $t$ for Gaussian input.
A key contribution of this experiment is the log-domain trapezoid integration method. Compared to standard trapezoid integration on the $t$-grid, the log-domain approach improved estimation accuracy by over an order of magnitude, reducing median relative errors from approximately 26% to 0.4%. This substantial improvement arises from: (1) uniform grid spacing in the log domain, (2) reduced convexity bias, and (3) better numerical handling of the wide dynamic range in $t$ values.
The tail correction, based on the asymptotic behavior $\text{mmse}(t)\sim\text{tr Cov}(X)$ for large $t$, proves essential for maintaining accuracy across the full integration range. Without tail correction, the 90th percentile error degrades significantly to over 40%.
These validation results on Gaussian inputs, where analytical solutions exist, provide strong evidence for the reliability of the proposed Score-to-Fisher bridge method and related theoretical framework presented in the previous sections.
VI-A5 Noise-conditional model
We investigate a noise-conditional score model $s_{\bm{\theta}}(\bm{y},t)$ (Scheme B) that takes both observation $\bm{y}$ and noise variance $t$ as inputs, eliminating the need for separate per-$t$ training. The model uses Gaussian Fourier projection to embed scalar $t$ values, which are injected into each hidden layer. Training samples $t$ from a log-uniform distribution over $Pt_{\min},Pt_{\max}]$ with loss weighting $\lambda(t)=t$ for 20,000 steps.
For Gaussian input with $n=4$ and $P=1.0$, this approach achieves 6.61% median relative error across 12 evaluation points spanning $t\in[0.005,50]$. Figure 4 demonstrates reasonable agreement with analytical values across four orders of magnitude in noise variance. We can observe the tendency that the conditional DSM estimate slightly underestimates the MI at low noise variance.
Figure 4: Noise-conditional model for mutual information estimation results for Gaussian input ($n=4$, $P=1$). Scheme B was used.
This unified framework requires only one neural network instead of 12 separate models, significantly reducing computational cost while eliminating the extreme outliers observed in per-$t$ training (Scheme A) and providing smooth interpolation across the entire noise range.
VI-B Discrete input case
VI-B1 Mutual Information Estimation
To validate the approach for discrete input distributions, we consider Binary Phase Shift Keying (BPSK) modulation with $X\in\{-\sqrt{P},+\sqrt{P}\}$ (equal probability), $P=1.0$, and dimension $n=1$. Unlike continuous distributions, BPSK admits exact mutual information calculation via numerical integration of the Gaussian mixture distribution providing ground truth for validation.
We implement per-$t$ training (Scheme A) over $M=12$ logarithmically spaced points in $t\in[0.005,50]$, with 1000 training iterations per noise level and 200,000 Monte Carlo samples for Fisher information estimation.
Figure 5 compares DSM estimates with exact BPSK mutual information and the Gaussian input baseline $I_{\text{Gauss}}(X;Y_{t})/n=\frac{1}{2}\log(1+P/t)$. The DSM estimates achieve excellent agreement with exact values, with median relative errors of 0.6% for low noise ($t<0.1$) and 2.8% for mid-range ($0.1\leq t<5$).
Figure 5: Mutual information estimation for BPSK input ($n=1$, $P=1$). Dotted line: Gaussian input theoretical baseline; solid circles: BPSK exact MI via numerical integration; dashed squares: DSM estimates.
VI-C Linear Gaussian channel case
To validate the method beyond the standard AWGN case, we consider the linear Gaussian channel:
$Y_{t}=AX+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (53)
where $A\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$ with $P=1.0$, $n=4$.
We train separate score networks (3 hidden layers, 128 units each) for $M=10$ logarithmically spaced noise levels in $t\in[P/200,50P]$. Each network is trained for 300 iterations with 8192-sample mini-batches, and Fisher information is estimated using 100,000 Monte Carlo samples per noise level.
Figure 6 compares the DSM-based estimates with the analytical solution (per symbol) $I(X;Y_{t})/n=(1/2)\log\det(I_{n}+(P/t)AA^{\top})$. The results demonstrate excellent agreement across 4 orders of magnitude in noise variance, confirming that the Score-to-Fisher bridge method generalizes successfully to linear Gaussian channels with nontrivial correlation structure.
Figure 6: Mutual information estimation for linear Gaussian channel $Y_{t}=AX+Z_{t}$ ($n=4$, $P=1$). Circles: analytical ground truth; triangles: DSM estimates.
VI-D Nonlinear channel case
To demonstrate the effectiveness of the proposed method for nonlinear channels where closed-form solutions are unavailable, we consider a composite channel with linear mixing followed by elementwise tanh nonlinearity:
$Y_{t}=\tanh(AX)+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (54)
where $A\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $\tanh(\cdot)$ is applied elementwise. This model represents practical scenarios where signals undergo linear transformations before nonlinear processing.
VI-D1 Experimental Setup
We use the same DSM configuration as in the linear case: $n=4$, $P=1.0$, with $M=12$ logarithmically spaced noise levels in $t\in[P/200,50P]$. The orthogonal matrix $A$ is generated via QR decomposition of a random Gaussian matrix. For each $t$, we train score networks for 400 iterations using 8192-sample mini-batches. Fisher information estimation uses 100,000 Monte Carlo samples per noise level.
VI-D2 KDE-LOO Baseline
Since analytical solutions are unavailable for nonlinear channels, we implement a Kernel Density Estimation with Leave-One-Out (KDE-LOO) baseline [6] that directly estimates mutual information from its definition. For samples $\{(\bm{w}_{j},\bm{y}_{j})\}_{j=1}^{N}$ with $\bm{y}_{j}=\bm{w}_{j}+\sqrt{t}\bm{\epsilon}_{j}$, the KDE-LOO estimator is defined by
$\hat{I}_{\text{KDE}}(X;Y_{t})=\frac{1}{N}\sum_{i=1}^{N}\Bigg[$ $-\frac{\|\bm{y}_{i}-\bm{w}_{i}\|^{2}}{2t}$
$-\log\left(\frac{1}{N-1}\sum_{j\neq i}e^{-\frac{\|\bm{y}_{i}-\bm{w}_{j}\|^{2}}{2t}}\right)\Bigg].$ (55)
This estimator is derived from the plug-in principle applied to $I(X;Y_{t})=\mathbb{E}[\log p_{Y_{t}|X}(\bm{y}|\bm{x})]-\mathbb{E}[\log p_{Y_{t}}(\bm{y})]$, where the marginal density $p_{Y_{t}}(\bm{y})$ is approximated using LOO kernel density estimation to avoid overfitting. We use $N=20,000$ samples with full-sum computation (no $k$-nearest neighbor approximation) to ensure high accuracy.
VI-D3 Results and Discussion
Figure 7 presents the comparison between DSM-based estimates and the KDE-LOO baseline for the composite nonlinear channel. The results demonstrate remarkable agreement between the two methods across the entire noise variance range, with both curves nearly overlapping.
Figure 7: Mutual information estimation for composite nonlinear channel $Y_{t}=\tanh(AX)+Z_{t}$ ($n=4$, $P=1$). Circles: KDE-LOO baseline; triangles: DSM estimates.
Key observations include:
  1. 1.
    Excellent agreement: The DSM estimates closely match the KDE-LOO baseline across four orders of magnitude in noise variance, with relative errors typically below 5%.
  2. 2.
    Computational efficiency: While KDE-LOO requires $O(N^{2})$ distance computations per noise level, DSM achieves comparable accuracy with $O(N)$ score evaluations after one-time training.
  3. 3.
    Nonlinear validation: This result provides strong evidence for the Score-to-Fisher bridge approach in scenarios where analytical benchmarks are unavailable.
The success of this experiment confirms that the proposed method generalizes effectively to handle composite channels combining linear mixing and nonlinear transformations, making it a practical tool for mutual information estimation in realistic communication systems.
VI-E Internal validation via Fisher information
To further validate the theoretical foundations of our approach, we conduct an internal consistency check by estimating Fisher information through two independent routes within the proposed framework. For the same composite nonlinear channel $Y_{t}=\tanh(AX)+Z_{t}$, we compare:
  1. 1.
    Score-based route: $J_{\text{score}}(Y_{t})\equiv\mathbb{E}[\|s_{\bm{\theta}}(Y_{t})\|^{2}]$ from trained score networks
  2. 2.
    MMSE-based route: $J_{\text{mmse}}(Y_{t})\equiv{(nt-\text{mmse}_{W}(t))}/{t^{2}}$ from trained denoisers
Both routes should yield identical Fisher information if the theoretical framework is sound.
Figure 8 presents the comparison between the two estimation routes across the noise variance range. At the low noise variance, the MMSE-based route underestimates the Fisher information due to the difficulty of the learning of the optimal denoiser at low noise variance. However, at the mid to high noise variance, the two routes show excellent agreement, with both curves nearly overlapping throughout most of the range. This internal consistency check provides strong evidence for the correctness of our theoretical approach and implementation.
Figure 8: Internal validation via Fisher information estimation. Circles: score-based route; triangles: MMSE-based route.
This validation experiment is particularly valuable as it provides a self-contained verification of the method’s theoretical consistency, independent of external baselines or ground truth comparisons.
VII Conclusion
We have presented a novel approach for mutual information estimation in nonlinear Gaussian noise channels via the Score-to-Fisher bridge. The method leverages DSM to estimate Fisher information, which is then converted to mutual information through the I-MMSE relation and log-domain integration with tail correction.
Experimental validation across linear Gaussian channels, composite nonlinear channels, and internal Fisher consistency checks demonstrates excellent agreement with analytical solutions and baseline methods. The approach achieves computational efficiency with $O(N)$ evaluations per noise level while maintaining high accuracy across wide dynamic ranges of the noise variance.
The Score-to-Fisher bridge methodology provides a practical and theoretically sound framework for mutual information estimation in realistic nonlinear Gaussian noise channel models where closed-form solutions are unavailable. Accurate estimation of mutual information plays a crucial role in assessing the fundamental limits of coded communication systems, and also serves as a valuable guideline for system design choices such as coding schemes, preprocessing strategies, and detection or decoding algorithms.
Acknowledgments
This work was supported by JST, CRONOS, Japan Grant Number JPMJCS25N5.

### v2

@@
 Abstract
-We present a numerical method to evaluate mutual information (MI) in nonlinear Gaussian noise channels by using denoisingscore matching learning for estimating the score function of channel output.
-Via de Bruijn’s identity and the I–MMSE relation, Fisher information estimated from the learned score yields accurate estimates of MI through an integral representation of MI for a variety of priors and channel nonlinearities (e.g., elementwise nonlinearity).
+We present a numerical method to evaluate mutual information (MI) in nonlinear Gaussian noise channels by using denoising score matching (DSM) learning for estimating the score function of channel output.
+Via de Bruijn’s identity, Fisher information estimated from the learned score function yields accurate estimates of MI through a Fisher integral representation for a variety of priors and channel nonlinearities.
 In this work, we propose a comprehensive theoretical foundation for the Score-to-Fisher bridge methodology, along with practical guidelines for its implementation.
 We also conduct extensive validation experiments, comparing our approach with closed-form solutions and a kernel density estimation baseline.
-The results of our numerical experiments demonstrate that the proposed method is both practical and efficient for mutual information estimation in nonlinear Gaussian noise channels.
+The results of our numerical experiments demonstrate that the proposed method is both practical and efficient for MI estimation in nonlinear Gaussian noise channels.
 I Introduction
@@
 $Y_{t}=f(X)+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (1)
-and study how to estimate the mutual information $I(X;Y_{t})$ efficiently and accurately.
-The nonlinear front-end can be represented by a deterministic function $f$.
+and study how to estimate the mutual information (MI) $I(X;Y_{t})$ efficiently and accurately.
+A nonlinear front-end function can be represented by a deterministic function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
 Such models arise widely due to nonlinear front-end saturation/clipping, DAC/ADC nonlinearity, photonic modulators, optical fibers, or learned encoders implemented with neural networks.
-The value of the mutual information $I(X;Y_{t})$ is important for evaluating the ultimate performance of coded communication systems and it provides useful information for designing the system such as coding, preprocessing, or decoding/detection algorithms.
+The value of the mutual information $I(X;Y_{t})$ is important for evaluating the ultimate performance of coded communication systems; it provides useful information for designing the system such as coding, preprocessing, or decoding/detection algorithms.
+In a sensing system, the nonlinear front-end can be interpreted as a details of sensing process.
+The value of the mutual information $I(X;Y_{t})$ indicates how much information is transferred from the sensing target to the output.
 When $f$ is nonlinear, evaluating $I(X;Y_{t})$ is, in general, challenging in practice because accurate posterior means $\mathbb{E}[X|Y_{t}]$ and normalized likelihoods are inaccessible or computationally intractable.
-For linear Gaussian channels, the de Bruijn identity and the I–MMSE relation yield
-$\frac{d}{dt}I(X;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}=-\frac{1}{2t^{2}}\mathrm{mmse}(t),$ (2)
-so that the mutual information $I(X;Y_{t})$ can be obtained by integrating either the Fisher information $J(Y_{t})$ or the MMSE.
-Our key observation is that the deterministic nonlinearity can be absorbed into a transformed input $W=f(X)$, so that
-$I(X;Y_{t})=I(W;Y_{t}),\quad Y_{t}=W+Z_{t},$
-holds for any deterministic function $f$. Hence, for general $f$ we can still rely on AWGN identities:
-$\mathrm{mmse}_{W}(t)=nt-t^{2}J(Y_{t}),$
-and consequently obtain $I(X;Y_{t})$ by integrating $J(Y_{t})$ or $\mathrm{mmse}_{W}(t)$.
-In this paper, we propose a score-based estimator of mutual information tailored to $Y_{t}=f(X)+Z_{t}$. The method learns the score of channel output, or the marginal $p_{Y_{t}}$,
-$s_{Y_{t}}(\bm{y})\equiv\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y}),$ (3)
-using denoising score matching (DSM) and then obtain an estimate of the Fisher information $\widehat{J}(Y_{t}).$ From this, we obtain an estimate of the mutual information via an integral representation of MI:
-$\widehat{I}(X;Y_{T})=\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-\widehat{J}(Y_{t})\Big)\,dt.$ (4)
-Crucially, the approach never requires access to the posterior $p_{X|Y}(\bm{x}|\bm{y})$.
-All we need are forward samples $\bm{x}\sim p_{X}$, evaluations of $\bm{w}=f(\bm{x})$, and Gaussian noise addition, which enables a single-loop estimator of MMSE and the Fisher information.
+Let $W=f(X)$ be a transformed input.
+The de Bruijn identity immediately yields
+$\frac{d}{dt}I(W;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}$ (2)
+so that the mutual information $I(X;Y_{t})$ can be obtained by integrating the Fisher information $J(Y_{t})$.
+Our key observation is that the Fisher information $J(Y_{t})$ can be estimated from the score function of $Y_{t}$ which can be estimated by denoising score matching (DSM) learning from samples of $Y_{t}$, and that the deterministic nonlinearity can be absorbed into a transformed input $W=f(X)$, so that
+$I(X;Y_{t})=I(W;Y_{t}),\quad Y_{t}=W+Z_{t},$ (3)
+holds for any deterministic function $f$. This equality is the basis of our score-based estimator of MI.
+By integrating (2) over $t\in[T,\infty)$, for general $f$, we have the Fisher integral representation of MI:
+$I(X;Y_{T})=\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-J(Y_{t})\Big)\,dt$ (4)
+that depends only on the Fisher information of $Y_{t}$.
+In this paper, we propose a score-based estimator of MI tailored to $Y_{t}=f(X)+Z_{t}$ [^1] A software implementation to reproduce the numerical results in this paper is available at https://github.com/wadayama/score-to-fisher-mi.
+The method learns the score of channel output, or the marginal density $p_{Y_{t}}$, using DSM learning and then obtain an estimate of the Fisher information $\widehat{J}(Y_{t}).$ From this, we obtain an estimate of the mutual information via an Fisher integral representation (4).
+Crucially, the approach never requires access to the posterior $p_{X|Y}(\cdot|\cdot)$.
+All we need is possibility of forward channel simulation, i.e., we assume that generation of samples $\bm{x}\sim p_{X}(\cdot)$, evaluations of $\bm{w}=f(\bm{x})$, and Gaussian noise addition are computationally feasible.
 The main contributions of this paper are summarized as follows:
-  • Score-only MI estimation for nonlinear front-ends.
-We formalize a pipeline— denoising score matching of $p_{Y_{t}}$ $\Rightarrow$ Fisher estimation $\Rightarrow$ MMSE/MI integration—that applies to arbitrary deterministic $f$ as soon as $p_{Y_{t}}$ can be sampled.
-  • Unified theoretical grounding.
-Leveraging de Bruijn and I–MMSE, we derive $\mathrm{mmse}_{W}(t)=nt-t^{2}J(Y_{t})$ and connect $I(X;Y_{t})$ to $I(W;Y_{t})$ with $W=f(X)$, extending linear-Gaussian identities to general nonlinear front-ends without ambiguity.
-  • Practical recipe.
-We detail implementation choices that make the estimator stable and reproducible: logarithmic grids in $t$, tail corrections for the MI integral, and a diagnostic “double estimate” that compares a Tweedie-based direct MMSE with the Fisher-based MMSE.
-  • Validation and use-cases. We verify the estimator on Gaussian inputs where closed forms are available, and report experiments on composite nonlinear channels.
-In linear Gaussian channels, I–MMSE and multivariate de Bruijn provide differential and integral representations of information quantities have been extensively studied.
+  • Score-only MI estimation for nonlinear front-ends: We formalize a pipeline— DSM learning of $p_{Y_{t}}$ $\Rightarrow$ Fisher estimation $\Rightarrow$ Fisher integral representation—that applies to arbitrary deterministic $f$ as soon as $p_{Y_{t}}$ can be easily sampled.
+  • Practical recipe: We detail implementation choices that make the estimator stable and reproducible: logarithmic grids in $t$, tail corrections for the MI integral.
+  • Validation and use-cases: We verify the estimator on Gaussian inputs where closed forms are available, and report experiments on composite nonlinear channels.
+In linear Gaussian channels, multivariate de Bruijn identity provides differential and integral representations of information quantities.
 For nonlinear $f$, however, posterior-free and normalization-free MI estimators remain limited.
-Building on score matching and denoising-based score learning, our approach attaches directly to the marginal $p_{Y_{t}}$, circumventing posterior computations and partition functions.
-The estimate of mutual information can be used many applications such as design of coding systems, optimization of preprocessing, and design of detection and decoding algorithms.
+Building on DSM learning, our approach attaches directly to the marginal density $p_{Y_{t}}$, circumventing posterior computations and partition functions.
+Moreover, comparing with the conventional kernel density estimation (KDE) based estimators, our approach is robust to the high-dimensionality of the input space.
+The estimate of MI can be used for many applications such as design of coding systems, optimization of preprocessing, sensor design, and system optimization, and design of detection and decoding algorithms.
 II Preliminaries
@@
 The (scalar) Fisher information is defined via the score function $s_{U}(\bm{u})\equiv\nabla_{\bm{u}}\log p_{U}(\bm{u})$ as
-$J(U)\equiv\mathbb{E}\!\left[\|\nabla_{\bm{u}}\log p_{U}(\bm{u})\|^{2}\right]\quad\text{with }\bm{u}\sim p_{U}$ (5)
+$J(U)\equiv\mathbb{E}_{\bm{u}\sim p_{U}(\cdot)}\!\left[\|\nabla_{\bm{u}}\log p_{U}(\bm{u})\|^{2}\right]$ (5)
 see, e.g., [1, 2].
@@
 Arnold et al.
-[3] proposed a simulation-based framework that computes upper and lower bounds on the information rate by combining auxiliary finite-state channel models with forward–backward recursions and importance sampling; the method is broadly applicable to ISI and finite-state channels where exact evaluation is infeasible.
+[3] proposed a simulation-based framework that computes upper and lower bounds on the information rate by combining auxiliary finite-state channel models with forward–backward recursions and importance sampling; the method is broadly applicable to intersymbol interference (ISI) and finite-state channels where exact evaluation is infeasible.
 Building on this line, Sadeghi et al.
 [4] optimized these bounds over the choice of auxiliary models, yielding systematically improved estimates and a practical tool for design and benchmarking of channels with memory.
-Another line of work estimates $I(X;Y)$ directly from its definition via simulation-based plug-in estimators: one approximates $p_{Y}$ (or $h(Y)$) using Kernel Density Estimation (KDE) or mixtures [5] or employs $k$-nearest neighbors (kkNN) estimators [6].
+Another line of work estimates the MI directly from its definition via simulation-based plug-in estimators: one approximates $p_{Y}$ (or $h(Y)$) using KDE [5] or employs $k$-nearest neighbors (kkNN) estimators [6].
 In the additive Gaussian setting, recent theory shows that plug-in estimation of $h(X+Z)$ is in fact optimal [7].
@@
 where $t$ represents the noise variance.
+The matrix $\mathbf{I}_{n}$ is the $n\times n$ identity matrix.
 Let $p_{Y_{t}}$ denote the marginal density of $Y_{t}$.
-The score function of $p_{Y_{t}}$ is defined by $s_{Y_{t}}(\bm{y})\equiv\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y}).$ Throughout, we assume that sampling from $p_{X}$ is computationally feasible by forward sampling or Langevin sampling using the score $\nabla_{\bm{x}}\log p_{X}(\bm{x})$ (e.g., Rossky et al.
+The score function of $p_{Y_{t}}$ is defined by
+$s_{Y_{t}}(\bm{y})\equiv\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y}).$ (7)
+Throughout, we assume that sampling from $p_{X}$ is computationally feasible by forward sampling or Langevin sampling using the score $\nabla_{\bm{x}}\log p_{X}(\bm{x})$ (e.g., Rossky et al.
 [8], Welling and Teh [9]).
 The Fisher information of $Y_{t}$ is given by
-$J(Y_{t})\equiv\mathbb{E}_{Y_{t}}\bigl[\|s_{Y_{t}}(\bm{y})\|^{2}\bigr].$ (7)
-Under standard regularity conditions, the following Tweedie identity [10, 11, 12] holds for $t>0$:
+$J(Y_{t})\equiv\mathbb{E}_{\bm{y}\sim p_{Y_{t}}(\cdot)}\bigl[\|s_{Y_{t}}(\bm{y})\|^{2}\bigr],$ (8)
+where $\|\cdot\|$ denotes the Euclidean norm. Under standard regularity conditions, the following Tweedie identity [10, 11, 12] holds for $t>0$:
 $\mathbb{E}[X|Y_{t}=\bm{y}]=\bm{y}+ts_{Y_{t}}(\bm{y}).$ (9)
@@
 II-E de Bruijn identity
-Let $h(Y_{t})$ denote the differential entropy of $Y_{t}$. Then the (multivariate) de Bruijn identity gives
+Let $h(Y_{t})$ denote the differential entropy of $Y_{t}$. Consider the AWGN channel setup defined by (6). Then the (multivariate) de Bruijn identity gives
 $\frac{d}{dt}h(Y_{t})=\frac{1}{2}J(Y_{t}).$ (14)
-The de Bruijn identity was discussed in [15, 16].
+The de Bruijn identity was discussed in [15, 16], which is proved by using the heat partial differential equation.
 II-F Denoising score matching (DSM) method
@@
 where this equivalence was shown by Vincent [18].
-This means that minimizing the DSM loss function is equivalent to minimizing the expectation of the squared distance between the learned score function and the true score function, which is sometimes called the Fisher divergence.
-This means that the unique minimizer of the DSM loss function satisfies $s_{\bm{\theta}}(\bm{y})\simeq\nabla_{\bm{y}}\log p_{\sigma}(\bm{y})$ in high probability.
+This means that minimizing the DSM loss function is equivalent to minimizing the expectation of the squared distance between the learned score function and the true score function, which is often called the Fisher divergence.
+This means that the unique minimizer of the DSM loss function satisfies $s_{\bm{\theta}}(\bm{y})\simeq\nabla_{\bm{y}}\log p_{\sigma}(\bm{y})$.
 Equivalently, since $\bm{v}-\bm{w}=\sigma\bm{\epsilon}$, one may implement the loss function as $(\bm{v}-\bm{w})/\sigma^{2}$:
@@
 which is an equivalent form of the DSM loss function.
-In practice, we use a neural network to parameterize the score model $s_{\bm{\theta}}(\bm{v})$.
+In practice, we use a neural network to parameterize the score model $s_{\bm{\theta}}(\cdot)$.
 Such a score model is also called a score network.
@@
 III Estimating Mutual Information in AWGN Channels
-In this section, we present a method to estimate the MMSE and MI for AWGN channels described by
+In this section, we present a method to estimate the MI for AWGN channels described by
 $Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),\quad t>0,$ (19)
 where $X\sim p_{X}$. The results obtained in this section will be extended to general nonlinear channels in Section V.
-III-A MMSE–Fisher relation
+III-A MI–Fisher relation
 For any Gaussian channel, the following proposition holds.
@@
 For all $t>0$, the identity
-$\mathrm{mmse}(t)=nt-t^{2}J(Y_{t}).$ (19)
+$\frac{d}{dt}I(X;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}$ (20)
 holds [13, 14, 21].
 Proof.
-A sketch of proof is the following. Using $I(X;Y_{t})=h(Y_{t})-h(Z_{t})$ with $h(Z_{t})=(n/2)\log(2\pi e\,t)$, we have
+A sketch of proof is the following. Using the definition of MI,
+$I(X;Y_{t})=h(Y_{t})-h(Z_{t})$ (21)
+with $h(Z_{t})=(n/2)\log(2\pi e\,t)$, we have
 $\frac{d}{dt}I(X;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}$ (22)
-by de Bruijn identity. Equating this with the I–MMSE form
-$\frac{d}{dt}I(X;Y_{t})=-\frac{1}{2t^{2}}\mathrm{mmse}(t)$ (21)
-gives $\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$. ∎
-Reeves et al.
-[21] investigated this identity in the context of linear Gaussian channels, highlighting its implications for non-Gaussian input distributions.
+by de Bruijn identity. ∎
+It can be remarked that, from the proposition, we immediately obtain the MMSE–Fisher relation
+$\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$ (23)
+by the I–MMSE relation.
+Note that Reeves et al.
+[21] investigated related identities in the context of linear Gaussian channels, highlighting its implications for non-Gaussian input distributions.
 Their work extends the I–MMSE framework beyond the classical Gaussian setting.
-III-B Integral representations of mutual information
+III-B Fisher Integral representations of mutual information
 From the identity (20) and the fact that $\lim_{t\to\infty}I(X;Y_{t})=0$ (since $Y_{t}$ and $X$ become independent in the limit), integrating over $t\in[T,\infty)$ yields
-$I(X;Y_{T})=$ $\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-J(Y_{t})\Big)\,dt$ (22)
-$=$ $\frac{1}{2}\int_{T}^{\infty}\!\frac{\mathrm{mmse}(t)}{t^{2}}\,dt.$ (23)
+$I(X;Y_{T})=$ $\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-J(Y_{t})\Big)\,dt.$ (24)
 The integrand decays as $O(1/t^{2})$, since
 $0\ \leq\ \frac{n}{t}-J(Y_{t})\;=\;\frac{\mathrm{mmse}(t)}{t^{2}}\ \leq\ \frac{\operatorname{tr}\operatorname{Cov}(X)}{t^{2}}.$ (25)
-The integral representation directly follows by combining the de Bruijn identity [15] with the I–MMSE relation [13, 14].
+The integral representation directly follows by combining the de Bruijn identity [15] with the I–MMSE relation [13, 14] as well.
 III-C Estimation of Fisher Information
-Using DSM method, we can estimate the score function of $Y_{t}$. With $p=p_{X}$ and $\sigma^{2}=t$, the learned score model coincides with the score of $Y_{t}=X+Z_{t}$:
-$s_{\bm{\theta}}(\cdot)\approx\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y})=s_{Y_{t}}(\bm{y}),$ (25)
+Using DSM learning, we can estimate the score function of $Y_{t}$. With $p\leftarrow p_{X}$ and $\sigma^{2}\leftarrow t$, the learned score model can approximate the score of $Y_{t}=X+Z_{t}$:
+$s_{\bm{\theta}}(\bm{y})\approx\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y})=s_{Y_{t}}(\bm{y}),$ (26)
 where $s_{\bm{\theta}}(\cdot)$ can be implemented with a neural network with parameter vector $\bm{\theta}$.
-This enables estimation of $J(Y_{t})=\mathbb{E}\|s_{Y_{t}}(Y_{t})\|^{2}$ from samples of $\bm{x}\sim p_{X}$ and Gaussian noises.
+This enables estimation of $J(Y_{t})=\mathbb{E}_{\bm{y}\sim p_{Y_{t}}(\cdot)}\|s_{Y_{t}}(\bm{y})\|^{2}$ from samples of $\bm{x}\sim p_{X}$ and Gaussian noises.
 In particular, for a fixed noise level $t$, the DSM loss function can be written as
 $\mathcal{L}_{t}(\bm{\theta})\;\equiv\;\mathbb{E}_{\bm{x},\bm{\epsilon}}\!\left[\left\|\,s_{\bm{\theta}}(\bm{x}+\sqrt{t}\,\bm{\epsilon})+\bm{\epsilon}/\sqrt{t}\,\right\|^{2}\right]$ (27)
-$=\mathbb{E}_{Y_{t}}\!\left[\left\|\,s_{\bm{\theta}}(Y_{t})-s_{Y_{t}}(Y_{t})\,\right\|^{2}\right]+\text{const},$ (27)
+$=\mathbb{E}_{\bm{y}}\!\left[\left\|\,s_{\bm{\theta}}(\bm{y})-s_{Y_{t}}(\bm{y})\,\right\|^{2}\right]+\text{const},$ (28)
 where the constant term does not depend on $\bm{\theta}$.
-Hence, with a sufficiently rich function class for $s_{\bm{\theta}}$ (e.g., a universal neural network) and sufficiently many samples, any sequence of empirical minimizers achieves $\mathbb{E}_{Y_{t}}\!\left[\left\|\,s_{\bm{\theta}}(Y_{t})-s_{Y_{t}}(Y_{t})\,\right\|^{2}\right]\rightarrow 0,$ i.e., the learned score converges to the true score in the Fisher divergence sense.
+Hence, with a sufficiently rich function class for $s_{\bm{\theta}}$ (e.g., a universal neural network) and sufficiently many samples, any sequence of empirical minimizers achieves $\mathbb{E}_{\bm{y}}\!\left[\left\|\,s_{\bm{\theta}}(\bm{y})-s_{Y_{t}}(\bm{y})\,\right\|^{2}\right]\rightarrow 0,$ i.e., the learned score converges to the true score in the Fisher divergence sense.
 Moreover, with i.i.d. draws $\bm{y}_{i}=\bm{x}_{i}+\sqrt{t}\,\bm{\epsilon}_{i}$ from $p_{Y_{t}}(i=1,2,...,N)$, the Monte Carlo estimator:
@@
 III-D Gaussian input case for validation
-In this subsection we provide closed-form expressions for the Gaussian-input case $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ in the AWGN channel family $Y_{t}=X+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$.
+In this subsection we provide closed-form expressions for the Gaussian-input case $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ in the AWGN channel family $Y_{t}=X+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$ where $P>0$ represents the signal power.
 It is well known [2] that $I(X;Y_{t})=(n/2)\log(1+P/t)$; see also [13] for the relation to MMSE.
@@
 and $J(Y_{t})=\mathbb{E}\|s_{Y_{t}}(Y_{t})\|^{2}=n/(P+t)$.
-The remaining identities follow by substituting $J(Y_{t})$ into $\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$ and by $h(X\mid Y_{t})=h(Y_{t})-I(X;Y_{t})$ with $h(Y_{t})=(n/2)\log\big((2\pi e)(P+t)\big).$
+The remaining identities follow by substituting $J(Y_{t})$ into $\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$ and by $h(X|Y_{t})=h(Y_{t})-I(X;Y_{t})$ with $h(Y_{t})=(n/2)\log\big((2\pi e)(P+t)\big).$
 III-E Linear Gaussian channel case
@@
 ∎
-IV Estimating MMSE and Mutual Information
-Using the identity between $\mathrm{mmse}(t)$ and the Fisher information, we can evaluate $I(X;Y_{T})$ by numerical integration.
-In this section, we present details of the method to estimate the MMSE and MI for the AWGN channel family.
-IV-A Score-based Estimation Method for MI and MMSE
+IV Estimation of Mutual Information
+Using the score-based estimation of the Fisher information, we can evaluate $I(X;Y_{T})$ by numerical integration.
+In this section, we present details of the method to estimate the MI for the AWGN channel family described by (6).
+IV-A Score-based Estimation Method for MI
 The proposed estimation algorithm is summarized in Algorithm 1. The core of the algorithm is the DSM learning for $p_{Y_{t}}$ and the estimation of the Fisher information from the score model.
-Algorithm 1 Score-to-Fisher bridge: Estimation Algorithm for MI and MMSE
-1: Prepare a logarithmically spaced grid over $t$ (details of the log grid are described in Subsection IV-C).
+Algorithm 1 Score-to-Fisher bridge: Estimation of MI
+1: Prepare a logarithmically spaced grid over $t$ (details of the log grid are described in Subsection IV-B).
 2: Sample generation: Generate a random mini-batch $\mathcal{B}$ consisting of multiple pairs of $(\bm{x},\bm{\epsilon})$ for mini-batch training with forward sampling of $X$: $\bm{x}\sim p_{X},\ \bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{n}).$
-3: Denoising score matching (DSM) learning for $p_{Y_{t}}$: train the score model parameters by minimizing the empirical DSM loss:
+3: DSM learning for $p_{Y_{t}}$: train the score model parameters by minimizing the empirical DSM loss:
 $\sum_{(\bm{x},\bm{\epsilon})\in\mathcal{B}}\big\|\,s_{\bm{\theta}}(\bm{x}+\sqrt{t}\,\bm{\epsilon})\;+\;\bm{\epsilon}/\sqrt{t}\,\big\|^{2}.$ (38)
-Any stochastic optimizer, such as Adam, can be applied to minimized the empirical loss function.
+Any stochastic optimizer, such as Adam, can be applied to minimize the empirical loss function.
 4: Estimating the Fisher information: for each $t$, generate $\bm{y}_{1},\ldots,\bm{y}_{N}$ (e.g., by forward sampling $\bm{y}_{i}=\bm{x}_{i}+\sqrt{t}\,\bm{\epsilon}_{i}$), and compute the estimated Fisher information:
-$\widehat{J}_{t}\equiv\frac{1}{N}\sum_{i=1}^{N}\big\|s_{\theta}(\bm{y}_{i})\big\|^{2},$ (38)
+$\widehat{J}_{t}=\frac{1}{N}\sum_{i=1}^{N}\big\|s_{\bm{\theta}}(\bm{y}_{i})\big\|^{2}$ (39)
 as an approximation of $J(Y_{t})$.
-5: Estimating the MMSE:
-$\widehat{\mathrm{mmse}}(t)=nt-t^{2}\widehat{J}_{t}.$ (39)
-6: Estimating the MI via the integral representation:
+5: Estimating the MI via the Fisher integral representation:
 $\widehat{I(X;Y_{T})}=\frac{1}{2}\int_{T}^{\infty}\Big(\frac{n}{t}-\widehat{J}_{t}\Big)\,dt,$ (40)
@@
     $\mathbb{E}_{t\sim\mathcal{T}}\,\lambda(t)\ \mathbb{E}_{\bm{x},\bm{\epsilon}}\,\big\|s_{\bm{\theta}}(\bm{x}+\sqrt{t}\,\bm{\epsilon},t)+\bm{\epsilon}/\sqrt{t}\big\|^{2},$ (41)
-    where $\mathcal{T}$ is a training distribution over $t$ (e.g., logarithmically spaced as described later) and $\lambda(t)$ is a weight (e.g., $\lambda(t)=t$).
+    where $\mathcal{T}$ is a training distribution over $t$ (e.g., logarithmically spaced as described later) and $\lambda(t)$ is a weight function, e.g., $\lambda(t)=t$.
 In the case of Scheme B, the score network has two inputs: the channel output $\bm{y}$ and the noise level $t$.
-In generative modeling frameworks for image generation, Scheme B is often used to train a noise-conditional score model.
-The training distribution $\mathcal{T}$ is logarithmically spaced as described in Subsection IV-C.
-The weight $\lambda(t)$ is a weight function that is used to weight the loss function.
-IV-B Validating the Fisher-based estimate
-Although Algorithm 1 presented in the previous subsection is an asymptotically consistent method to estimate the MMSE and MI, the accuracy of the estimate is not clear in the case of finite samples.
-One possible source of the error is the estimation error of the Fisher information due to inaccurate estimation of the score model.
-Although we can expect the accurate score matching in the Fisher divergence sense in the case of infinite samples, the convergence rate depends on the prior distribution of the source.
-We also have to consider the numerical error of the numerical integration.
-In this section, we present a method to validate the estimated MMSE and MI. As a diagnostic purpose, we can compute the direct MMSE estimate based on the Tweedie formula:
-$\widehat{\mathrm{mmse}}_{\rm dir}(t)\equiv\frac{1}{N}\sum_{i=1}^{N}\big\|\bm{x}_{i}-\{\bm{y}_{i}(t)+t\,s_{\theta}(\bm{y}_{i}(t))\}\big\|^{2},$ (42)
-and check agreement with the Fisher-based estimate:
-$\widehat{\mathrm{mmse}}_{F}(t)\equiv nt-t^{2}\widehat{J}_{t}.$ (43)
-This consistency check is useful to validate the accuracy of the Fisher-based estimate.
-IV-C Log-domain trapezoid integration
-The integrand in the integration representation of MI is smooth and may have a long tail.
+IV-B Log-domain trapezoid integration
+The integrand in the Fisher integral representation (4) of the MI is smooth and may have a long tail.
 The choice of the integration grid is critical for achieving high numerical accuracy.
@@
 Rather than applying the trapezoid rule directly on the non-uniform $t$-grid, we perform the change of variables $u=\log t$ to obtain uniform spacing in the log domain.
-Using the integral representation
-$I(X;Y_{T})=\frac{1}{2}\int_{T}^{\infty}\!\left(\frac{n}{t}-J(Y_{t})\right)\,dt,$ (45)
-and substituting $u=\log t$ (so $dt=e^{u}du$ and $t=e^{u}$), we obtain
+Using the integral representation (4), and substituting $u=\log t$ (so $dt=e^{u}du$ and $t=e^{u}$), we obtain
 $I(X;Y_{T})=\frac{1}{2}\int_{\log T}^{\infty}\!\Big(n-e^{u}\,J(Y_{e^{u}})\Big)\,du.$ (43)
-Since the $t$-grid (44) corresponds to uniform spacing in $u=\log t$ with step size $\Delta_{u}\equiv\log(t_{\max}/t_{\min})/(M-1)$, we apply the standard trapezoid rule (uniform spacing):
+Since the $t$-grid (42) corresponds to uniform spacing in $u=\log t$ with step size $\Delta_{u}\equiv\log(t_{\max}/t_{\min})/(M-1)$, we apply the standard trapezoid rule with uniform spacing:
 $\widehat{I}(X;Y_{T})\approx\frac{1}{2}\sum_{k=k_{T}}^{M-2}\frac{\ell(u_{k})+\ell(u_{k+1})}{2}\,\Delta_{u}\;+\;\text{tail},$ (44)
-where $\ell(u)\equiv n-e^{u}\,\widehat{J}(Y_{e^{u}})$ (equivalently $\ell(t)=n-t\,\widehat{J}(Y_{t})$), and $k_{T}$ is the smallest index with $t_{k_{T}}\geq T$.
+where $\ell(u)\equiv n-e^{u}\,\widehat{J}(Y_{e^{u}})$, and $k_{T}$ is the smallest index with $t_{k_{T}}\geq T$.
 This log-domain integration offers several benefits over direct $t$-domain trapezoid integration: (1) uniform grid spacing reduces discretization error, (2) the transformation naturally handles a wide dynamic range in $t$, and (3) convexity bias is significantly reduced.
@@
 $Y_{t}=f(X)+Z_{t},$
-where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a deterministic function.
+where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a deterministic function [^2] An extension to a general case $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is straightforward but we assume $m=n$ for simplicity..
 The mutual information equivalence presented in the following proposition is a key observation which enables the score-based estimation to be applied to general nonlinear channels.
@@
 Then the conditional law of $Y_{t}$ given $X$ depends on $X$ only through $W$.
-This means that $X\to W\to Y_{t}$ forms a Markov chain and $Y_{t}\perp X\mid W$.
+This means that $X\to W\to Y_{t}$ forms a Markov chain and equivalently $Y_{t}\perp X\mid W$.
 We can prove mutual information equivalence in the following proposition.
@@
 V-B Score-based estimation method for nonlinear channels
-In the following discussin, we assume that the evalution of $f$ is computationally feasible. Namely, sampling from $p_{W}$ is feasible. Some parts of Algorithm 1 are replaced by the following steps:
-  • Sample $W$ by $W=f(X)$, where $X\sim p_{X}$.
-  • Estimate the Fisher information of $Y_{t}$ from the samples of $W$ by the score-based estimation method in Algorithm 1.
-  • Use the estimated Fisher information to evaluate the integral form of MI.
-  • Use the relationship $I(X;Y_{t})=I(W;Y_{t})$ to obtain the estimate of $I(X;Y_{t})$.
-The estimation procedure is still a single-loop procedure, which does not require the access to the posterior $p_{X|Y_{t}}$.
+In the following discussin, we assume that the evalution of $f$ is computationally feasible.
+Namely, sampling from $p_{W}$ is feasible.
+Some parts of Algorithm 1 should be replaced by the following steps:
+  • Sample $W$ by $W=f(X)$ where $X\sim p_{X}$.
+  • Estimate the Fisher information of $Y_{t}$ from the samples of $W$ by the score-based estimation in Algorithm 1.
+  • Use the estimated Fisher information to evaluate the Fisher integral form of MI.
+  • Use the equality $I(X;Y_{t})=I(W;Y_{t})$ to obtain the estimate of $I(X;Y_{t})$.
+The estimation procedure is still a single-loop procedure, which does not require the access to the posterior density $p_{X|Y_{t}}(\cdot|\cdot)$.
 VI Numerical experiments
@@
 VI-A1 Experimental Setup
-We implemented the DSM method with the following configuration: The score network is implemented as a fully connected neural network consisting of 3 hidden layers, each containing 128 units with SiLU (Swish) activation functions.
+We implemented the DSM learning for $p_{Y_{t}}$ with the following configuration: The score network is implemented as a fully connected neural network consisting of 3 hidden layers, each containing 128 units with SiLU (Swish) activation functions.
 The input and output dimensions match the problem dimension $n$, where we test with $n\in\{4,8,16\}$ to evaluate scalability.
@@
 Figure 1: Fisher information $J(Y_{t})$ versus noise variance $t$ for Gaussian input with $P=1$ and dimensions $n\in\{4,8,16\}$.
-Solid lines with circular markers show ground truth values $J(Y_{t})=n/(P+t)$, while dashed lines with triangular markers show DSM estimates $\hat{J}(Y_{t})$.
+Solid lines with circular markers show ground truth values $J(Y_{t})=n/(P+t)$, while dashed lines with triangular markers show estimated values $\hat{J}(Y_{t})$.
 The results confirm that:
@@
     The median relative error remains below 1.3% for all tested dimensions, validating the theoretical framework.
-VI-A3 Validating the Fisher-based MMSE Estimate
-To validate our Fisher-based MMSE estimation, we compare three methods using the same experimental setup with $n=4$: (1) true MMSE (analytical ground truth), (2) Tweedie-based direct estimation using
-$\widehat{\text{mmse}}_{\text{dir}}(t)\equiv\frac{1}{N}\sum_{i=1}^{N}\|\bm{x}_{i}-(\bm{y}_{i}+ts_{\bm{\theta}}(\bm{y}_{i}))\|^{2},$ (51)
-and (3) our Fisher-based method using
-$\widehat{\text{mmse}}_{\text{F}}(t)=nt-t^{2}\hat{J}(Y_{t}).$ (52)
-The noise variance range is set to $t\in[P/200,50P]$ to avoid numerical instabilities at extremely high noise levels.
-The results are shown in Figure 2.
-The two estimates are in excellent agreement with the true MMSE in the wide range of noise variance.
-Figure 2: Comparison of MMSE estimation methods for Gaussian input ($n=4$, $P=1$).
-VI-A4 Mutual Information Estimation
-We demonstrate the complete pipeline from the Score-to-Fisher bridge for mutual information estimation using the integral representation of MI.
+VI-A3 Mutual Information Estimation
+We demonstrate the complete pipeline from the Score-to-Fisher bridge for MI estimation using the integral representation of MI.
 The experiment uses the same configuration as previous tests with $n=4$ and $P=1.0$, implementing log-domain trapezoid integration with tail correction.
@@
 These validation results on Gaussian inputs, where analytical solutions exist, provide strong evidence for the reliability of the proposed Score-to-Fisher bridge method and related theoretical framework presented in the previous sections.
-VI-A5 Noise-conditional model
+VI-A4 Noise-conditional model
 We investigate a noise-conditional score model $s_{\bm{\theta}}(\bm{y},t)$ (Scheme B) that takes both observation $\bm{y}$ and noise variance $t$ as inputs, eliminating the need for separate per-$t$ training.
@@
 Figure 3 demonstrates reasonable agreement with analytical values across four orders of magnitude in noise variance.
-We can observe the tendency that the conditional DSM estimate slightly underestimates the MI at low noise variance.
+We can observe the tendency that the noise-conditional DSM estimate slightly underestimates the MI at low noise variance.
+This result may be due to insufficient training iterations for a conditional model.
+However, this unified framework requires only one neural network instead of 12 separate models, significantly reducing computational cost.
 Figure 3: Noise-conditional model for mutual information estimation results for Gaussian input ($n=4$, $P=1$). Scheme B was used.
-This unified framework requires only one neural network instead of 12 separate models, significantly reducing computational cost while eliminating the extreme outliers observed in per-$t$ training (Scheme A) and providing smooth interpolation across the entire noise range.
 VI-B Discrete input case
@@
 To validate the method beyond the standard AWGN case, we consider the linear Gaussian channel:
-$Y_{t}=AX+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (53)
-where $A\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$ with $P=1.0$, $n=4$.
+$Y_{t}={\bm{A}}X+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (48)
+where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$ with $P=1.0$, $n=4$.
 We train separate score networks (3 hidden layers, 128 units each) for $M=10$ logarithmically spaced noise levels in $t\in[P/200,50P]$.
@@
 The results demonstrate excellent agreement across 4 orders of magnitude in noise variance, confirming that the Score-to-Fisher bridge method generalizes successfully to linear Gaussian channels with nontrivial correlation structure.
-Figure 6: Mutual information estimation for linear Gaussian channel $Y_{t}=AX+Z_{t}$ ($n=4$, $P=1$). Circles: analytical ground truth; triangles: DSM estimates.
+Figure 5: Mutual information estimation for linear Gaussian channel $Y_{t}={\bm{A}}X+Z_{t}$ ($n=4$, $P=1$). Circles: analytical ground truth; triangles: DSM estimates.
 VI-D Nonlinear channel case
 To demonstrate the effectiveness of the proposed method for nonlinear channels where closed-form solutions are unavailable, we consider a composite channel with linear mixing followed by elementwise tanh nonlinearity:
-$Y_{t}=\tanh(AX)+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (54)
-where $A\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $\tanh(\cdot)$ is applied elementwise.
+$Y_{t}=\tanh({\bm{A}}X)+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (49)
+where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $\tanh(\cdot)$ is applied elementwise.
 This model represents practical scenarios where signals undergo linear transformations before nonlinear processing.
@@
 We use the same DSM configuration as in the linear case: $n=4$, $P=1.0$, with $M=12$ logarithmically spaced noise levels in $t\in[P/200,50P]$.
-The orthogonal matrix $A$ is generated via QR decomposition of a random Gaussian matrix.
+The orthogonal matrix $\bm{A}$} is generated via QR decomposition of a random Gaussian matrix.
 For each $t$, we train score networks for 400 iterations using 8192-sample mini-batches.
@@
 The results demonstrate remarkable agreement between the two methods across the entire noise variance range, with both curves nearly overlapping.
-Figure 7: Mutual information estimation for composite nonlinear channel $Y_{t}=\tanh(AX)+Z_{t}$ ($n=4$, $P=1$). Circles: KDE-LOO baseline; triangles: DSM estimates.
+Figure 6: Mutual information estimation for composite nonlinear channel $Y_{t}=\tanh({\bm{A}}X)+Z_{t}$ ($n=4$, $P=1$). Circles: KDE-LOO baseline; triangles: DSM estimates.
 Key observations include:
@@
   2. 2.
-    Computational efficiency: While KDE-LOO requires $O(N^{2})$ distance computations per noise level, DSM achieves comparable accuracy with $O(N)$ score evaluations after one-time training.
+    Computational efficiency: While KDE-LOO requires $O(N^{2})$ distance computations per noise level, DSM achieves comparable accuracy with $O(N)$ score evaluations after one-time training where $N$ is the number of samples.
   3. 3.
@@
 The success of this experiment confirms that the proposed method generalizes effectively to handle composite channels combining linear mixing and nonlinear transformations, making it a practical tool for mutual information estimation in realistic communication systems.
-VI-E Internal validation via Fisher information
-To further validate the theoretical foundations of our approach, we conduct an internal consistency check by estimating Fisher information through two independent routes within the proposed framework.
-For the same composite nonlinear channel $Y_{t}=\tanh(AX)+Z_{t}$, we compare:
-  1. 1.
-    Score-based route: $J_{\text{score}}(Y_{t})\equiv\mathbb{E}[\|s_{\bm{\theta}}(Y_{t})\|^{2}]$ from trained score networks
-  2. 2.
-    MMSE-based route: $J_{\text{mmse}}(Y_{t})\equiv{(nt-\text{mmse}_{W}(t))}/{t^{2}}$ from trained denoisers
-Both routes should yield identical Fisher information if the theoretical framework is sound.
-Figure 8 presents the comparison between the two estimation routes across the noise variance range.
-At the low noise variance, the MMSE-based route underestimates the Fisher information due to the difficulty of the learning of the optimal denoiser at low noise variance.
-However, at the mid to high noise variance, the two routes show excellent agreement, with both curves nearly overlapping throughout most of the range.
-This internal consistency check provides strong evidence for the correctness of our theoretical approach and implementation.
-Figure 8: Internal validation via Fisher information estimation. Circles: score-based route; triangles: MMSE-based route.
-This validation experiment is particularly valuable as it provides a self-contained verification of the method’s theoretical consistency, independent of external baselines or ground truth comparisons.
 VII Conclusion
-We have presented a novel approach for mutual information estimation in nonlinear Gaussian noise channels via the Score-to-Fisher bridge.
-The method leverages DSM to estimate Fisher information, which is then converted to mutual information through the I-MMSE relation and log-domain integration with tail correction.
-Experimental validation across linear Gaussian channels, composite nonlinear channels, and internal Fisher consistency checks demonstrates excellent agreement with analytical solutions and baseline methods.
+We have presented a novel approach for mutual information estimation in nonlinear Gaussian noise channels via the Score-to-Fisher bridge approach.
+The method leverages DSM learning of score model for $p_{Y_{t}}$ to estimate Fisher information, which is then converted to mutual information via the Fisher integral representation.
+Experimental validation across linear Gaussian channels, composite nonlinear channels, demonstrates excellent agreement with analytical solutions and baseline methods.
 The approach achieves computational efficiency with $O(N)$ evaluations per noise level while maintaining high accuracy across wide dynamic ranges of the noise variance.
-The Score-to-Fisher bridge methodology provides a practical and theoretically sound framework for mutual information estimation in realistic nonlinear Gaussian noise channel models where closed-form solutions are unavailable.
-Accurate estimation of mutual information plays a crucial role in assessing the fundamental limits of coded communication systems, and also serves as a valuable guideline for system design choices such as coding schemes, preprocessing strategies, and detection or decoding algorithms.
+The curse of dimensionality can be avoided due to the efficiency of the score model evaluation.
+The Score-to-Fisher bridge methodology provides a practical and theoretically sound framework for MI estimation in realistic nonlinear Gaussian noise channel models where closed-form solutions are unavailable.
+Accurate estimation of MI plays a crucial role in assessing the fundamental limits of coded communication systems, and also serves as a valuable guideline for system design choices such as coding schemes, preprocessing strategies, and detection or decoding algorithms.
 Acknowledgments

### v3

@@
 The results of our numerical experiments demonstrate that the proposed method is both practical and efficient for MI estimation in nonlinear Gaussian noise channels.
+Additionally, we discuss the theoretical connections between our score-based framework and thermodynamic concepts, such as partition function estimation and optimal transport.
 I Introduction
 We consider additive Gaussian noise channels with a deterministic nonlinear front-end,
-$Y_{t}=f(X)+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (1)
+$Y_{t}=f(X)+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n}),$ (1)
 and study how to estimate the mutual information (MI) $I(X;Y_{t})$ efficiently and accurately.
@@
 $\frac{d}{dt}I(W;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}$ (2)
-so that the mutual information $I(X;Y_{t})$ can be obtained by integrating the Fisher information $J(Y_{t})$.
+so that the mutual information $I(W;Y_{t})$ can be obtained by integrating the Fisher information $J(Y_{t})$.
 Our key observation is that the Fisher information $J(Y_{t})$ can be estimated from the score function of $Y_{t}$ which can be estimated by denoising score matching (DSM) learning from samples of $Y_{t}$, and that the deterministic nonlinearity can be absorbed into a transformed input $W=f(X)$, so that
@@
 that depends only on the Fisher information of $Y_{t}$.
-In this paper, we propose a score-based estimator of MI tailored to $Y_{t}=f(X)+Z_{t}$ [^1] A software implementation to reproduce the numerical results in this paper is available at https://github.com/wadayama/score-to-fisher-mi.
+In this paper, we propose a score-based estimator of MI [^1] A software implementation to reproduce the numerical results in this paper is available at https://github.com/wadayama/score-to-fisher-mi.
+tailored to the channel represented by $Y_{t}=f(X)+Z_{t}$.
 The method learns the score of channel output, or the marginal density $p_{Y_{t}}$, using DSM learning and then obtain an estimate of the Fisher information $\widehat{J}(Y_{t}).$ From this, we obtain an estimate of the mutual information via an Fisher integral representation (4).
@@
 All we need is possibility of forward channel simulation, i.e., we assume that generation of samples $\bm{x}\sim p_{X}(\cdot)$, evaluations of $\bm{w}=f(\bm{x})$, and Gaussian noise addition are computationally feasible.
-The main contributions of this paper are summarized as follows:
-  • Score-only MI estimation for nonlinear front-ends: We formalize a pipeline— DSM learning of $p_{Y_{t}}$ $\Rightarrow$ Fisher estimation $\Rightarrow$ Fisher integral representation—that applies to arbitrary deterministic $f$ as soon as $p_{Y_{t}}$ can be easily sampled.
-  • Practical recipe: We detail implementation choices that make the estimator stable and reproducible: logarithmic grids in $t$, tail corrections for the MI integral.
-  • Validation and use-cases: We verify the estimator on Gaussian inputs where closed forms are available, and report experiments on composite nonlinear channels.
+The main contributions of this paper are summarized as follows: (1)Score-only MI estimation for nonlinear front-ends: we formalize a pipeline: DSM learning of $p_{Y_{t}}$ $\Rightarrow$ Fisher estimation $\Rightarrow$ Fisher integral representation—that applies to arbitrary deterministic $f$ as long as $p_{Y_{t}}$ can be easily sampled.
+(2)Practical recipe: we detail implementation choices that make the estimator stable and reproducible: logarithmic grids in $t$, tail corrections for the MI integral.
+(3)Validation and use-cases: we verify the estimator on Gaussian inputs where closed forms are available, and report experiments on composite nonlinear channels.
 In linear Gaussian channels, multivariate de Bruijn identity provides differential and integral representations of information quantities.
@@
 The estimate of MI can be used for many applications such as design of coding systems, optimization of preprocessing, sensor design, and system optimization, and design of detection and decoding algorithms.
+In modern score-based modeling, the score function model serves as a tractable surrogate for otherwise intractable densities.
+Our score-to-fisher bridge leverages this view to turn mutual information into a one-dimensional integral of score-based functionals along Gaussian smoothing, enabling accurate numerical MI estimation without high-dimensional density estimation.
 II Preliminaries
@@
 The (scalar) Fisher information is defined via the score function $s_{U}(\bm{u})\equiv\nabla_{\bm{u}}\log p_{U}(\bm{u})$ as
-$J(U)\equiv\mathbb{E}_{\bm{u}\sim p_{U}(\cdot)}\!\left[\|\nabla_{\bm{u}}\log p_{U}(\bm{u})\|^{2}\right]$ (5)
+$J(U)\equiv\mathbb{E}_{\bm{u}\sim p_{U}(\cdot)}\!\left[\|\nabla_{\bm{u}}\log p_{U}(\bm{u})\|^{2}\right],$ (5)
 see, e.g., [1, 2].
 II-B Related works
+II-B1 Estimation of Mutual Information
 For channels with memory, simulation-based methods have been developed to evaluate information rates without closed-form expressions.
@@
 In the additive Gaussian setting, recent theory shows that plug-in estimation of $h(X+Z)$ is in fact optimal [7].
+II-B2 Theoretical Foundations and Thermodynamic Connections
+The integral representation used in this work is deeply rooted in the interplay between information theory and thermodynamics, particularly through the heat equation.
+This connection dates back to Stam [15], who derived the de Bruijn identity to prove the entropy power inequality, explicitly treating noise addition as a diffusion process.
+This line of inquiry was further developed by Brown [16] and Barron [17], who established the entropic central limit theorem, proving that the Fisher information and entropy monotonically converge to those of a Gaussian distribution during diffusion.
+In a modern context, Guo, Shamai, and Verdú [13] unified these concepts through the I-MMSE relationship, revealing that the derivative of mutual information is equivalent to the minimum mean-square error (and thus Fisher information) in Gaussian channels.
+Our proposed Score-to-Fisher bridge can be viewed as a data-driven operationalization of these classical theoretical frameworks: while Stam and Barron utilized these relations for proofs of inequalities, we leverage them for the exact numerical estimation of information quantities via learned score functions.
+II-B3 Diffusion Process and Fisher Information
+Recently, score-based representations have found applications beyond generative modeling, extending into information-theoretic analysis.
+For instance, Premkumar [23] introduced “neural entropy” to quantify the information stored in diffusion models, relating the total entropy production along a diffusion path to a time integral of a Fisher-divergence functional (i.e., an expected squared difference of score functions).
+While their framework is motivated by a non-equilibrium thermodynamic view of diffusion generative models (and the resulting quantity depends on the chosen forward process), it shares a fundamental mathematical connection with our approach: both convert global information measures into one-dimensional integrals of score-based functionals, through diffusion/heat-equation identities (de Bruijn’s identity in our additive Gaussian-noise setting).
 II-C AWGN channel and score function
 Let $X\in\mathbb{R}^{n}$ be a channel input random vector with known prior probability density function (PDF) $p_{X}$, and consider the AWGN channel family:
-$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),\quad t>0,$ (6)
+$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n}),\quad t>0,$ (6)
 where $t$ represents the noise variance.
-The matrix $\mathbf{I}_{n}$ is the $n\times n$ identity matrix.
+The matrix $\bm{I}_{n}$ is the $n\times n$ identity matrix.
 Let $p_{Y_{t}}$ denote the marginal density of $Y_{t}$.
@@
 $J(Y_{t})\equiv\mathbb{E}_{\bm{y}\sim p_{Y_{t}}(\cdot)}\bigl[\|s_{Y_{t}}(\bm{y})\|^{2}\bigr],$ (8)
-where $\|\cdot\|$ denotes the Euclidean norm. Under standard regularity conditions, the following Tweedie identity [10, 11, 12] holds for $t>0$:
+where $\|\cdot\|$ denotes the Euclidean norm. Under standard regularity conditions, the following Robbins-Tweedie identity [10, 11, 12] holds for $t>0$:
 $\mathbb{E}[X|Y_{t}=\bm{y}]=\bm{y}+ts_{Y_{t}}(\bm{y}).$ (9)
@@
 For the SNR-parameterized AWGN model
-$Y_{\gamma}=\sqrt{\gamma}X+N,\quad N\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n}),\quad N\perp X,$ (10)
+$Y_{\gamma}=\sqrt{\gamma}X+N,\quad N\sim\mathcal{N}(\bm{0},\bm{I}_{n}),\quad N\perp X,$ (10)
 the I–MMSE relation [13, 14] states that the identity
@@
 The de Bruijn identity was discussed in [15, 17], which is proved by using the heat partial differential equation.
+Intuitively, diffusion acts as a smoothing process that destroys information.
+As illustrated in Fig.
+1, this process is characterized by the decay of Fisher information $J(Y_{t})$ (which measures the ‘sharpness’ or structural complexity of the distribution) and the simultaneous increase in entropy $h(Y_{t})$.
+Our method computes the total information content by integrating this Fisher information profile.
+Figure 1: Visualization of the heat diffusion process for a 1D multimodal distribution (mixture of Gaussians).
+As time $t$ progresses, the distribution $p(x,t)$ becomes smoother, leading to an increase in differential entropy $h(X_{t})$ and a rapid decay in Fisher information $J(X_{t})$.
+The Score-to-Fisher Bridge method leverages this relationship via de Bruijn’s identity: the mutual information is essentially accumulated Fisher-information gap relative to the Gaussian baseline (the complexity of the distribution) over the diffusion time.
 II-F Denoising score matching (DSM) method
@@
 Let $p$ be a target density on $\mathbb{R}^{n}$ and define a Gaussian-smoothed sample:
-$\bm{v}=\bm{w}+\sigma\bm{\epsilon},\quad\bm{w}\sim p,\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n}).$ (15)
+$\bm{v}=\bm{w}+\sigma\bm{\epsilon},\quad\bm{w}\sim p,\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{n}).$ (15)
 Writing $p_{\sigma}\equiv p*\phi_{\sigma^{2}}$ for the convolution with a zero-mean Gaussian PDF $\phi_{\sigma^{2}}$ with variance $\sigma^{2}$, the DSM loss function for a parametric score model $s_{\bm{\theta}}:\mathbb{R}^{n}\!\to\!\mathbb{R}^{n}$ ($\bm{\theta}$ represents the vector of adjustable parameters) is given by
@@
 In this section, we present a method to estimate the MI for AWGN channels described by
-$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),\quad t>0,$ (19)
+$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n}),\quad t>0,$ (19)
 where $X\sim p_{X}$. The results obtained in this section will be extended to general nonlinear channels in Section V.
@@
 holds [13, 14, 22].
-Proof.
+Proof:
 A sketch of proof is the following. Using the definition of MI,
@@
 The integral representation directly follows by combining the de Bruijn identity [15] with the I–MMSE relation [13, 14] as well.
+Remark 1 (Connection to GEXIT function).
+The integrand in our estimation formula corresponds to the channel GEXIT function $g(t)$. Although typically defined via conditional entropy, it satisfies the following identities in the AWGN setting:
+$g(t)\equiv\frac{d}{dt}h(X|Y_{t})=-\frac{d}{dt}I(X;Y_{t})=\frac{n}{2t}-\frac{1}{2}J(Y_{t}).$ (26)
+By the area theorem, the integral of $g(t)$ characterizes the mutual information, thus linking our score-based Fisher integral to the GEXIT analysis commonly used in coding theory.
 III-C Estimation of Fisher Information
@@
 Moreover, with i.i.d. draws $\bm{y}_{i}=\bm{x}_{i}+\sqrt{t}\,\bm{\epsilon}_{i}$ from $p_{Y_{t}}(i=1,2,...,N)$, the Monte Carlo estimator:
-$\widehat{J}_{t}\triangleq\frac{1}{N}\sum_{i=1}^{N}\|s_{Y_{t}}(\bm{y}_{i})\|^{2}$ (29)
+$\widehat{J}_{t}\equiv\frac{1}{N}\sum_{i=1}^{N}\|s_{\bm{\theta}}(\bm{y}_{i})\|^{2}$ (30)
 satisfies $\widehat{J}_{t}\xrightarrow{\text{a.s.}}J(Y_{t})$ as $N\to\infty$ (law of large numbers), and $\sqrt{N}\,(\widehat{J}_{t}-J(Y_{t}))\Rightarrow\mathcal{N}\big(0,\operatorname{Var}(\|s_{Y_{t}}(\bm{y})\|^{2})\big)$ (central limit theorem), so the estimation error decays at rate $O(N^{-1/2})$.
@@
 III-D Gaussian input case for validation
-In this subsection we provide closed-form expressions for the Gaussian-input case $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ in the AWGN channel family $Y_{t}=X+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$ where $P>0$ represents the signal power.
+In this subsection we provide closed-form expressions for the Gaussian-input case $X\sim\mathcal{N}(\bm{0},P\,\bm{I}_{n})$ in the AWGN channel family $Y_{t}=X+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\bm{0},t\,\bm{I}_{n})$ where $P>0$ represents the signal power.
 It is well known [2] that $I(X;Y_{t})=(n/2)\log(1+P/t)$; see also [13] for the relation to MMSE.
 Proposition 2 (Closed forms for Gaussian inputs).
-Let $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ and $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$, independent. Then, for all $t>0$, we have the following closed forms:
+Let $X\sim\mathcal{N}(\bm{0},P\,\bm{I}_{n})$ and $Z_{t}\sim\mathcal{N}(\bm{0},t\,\bm{I}_{n})$, independent. Then, for all $t>0$, we have the following closed forms:
 $\frac{1}{n}I(X;Y_{t})=\frac{1}{2}\log\!\Big(1+\frac{P}{t}\Big),$ (31)
@@
 $\mathrm{mmse}(t)=n\,\frac{Pt}{P+t},\quad J(Y_{t})=\frac{n}{P+t}.$ (33)
-Sketch of proof: .
-From the Gaussian model $Y_{t}=X+Z_{t}$, it is known that $I(X;Y_{t})=(n/2)\log(1+P/t)$ holds. Moreover $Y_{t}\sim\mathcal{N}(\mathbf{0},(P+t)\mathbf{I}_{n})$, hence we have
+Proof:
+From the Gaussian model $Y_{t}=X+Z_{t}$, it is known that $I(X;Y_{t})=(n/2)\log(1+P/t)$ holds. Moreover $Y_{t}\sim\mathcal{N}(\bm{0},(P+t)\bm{I}_{n})$, hence we have
 $s_{Y_{t}}(\bm{y})=\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y})=-(P+t)^{-1}\bm{y}$ (34)
@@
 Consider a linear Gaussian channel model:
-$Y_{t}=AX+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (34)
-where $A\in\mathbb{R}^{n\times n}$ is a square, invertible matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$.
+$Y_{t}=\bm{A}X+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n}),$ (35)
+where $\bm{A}\in\mathbb{R}^{n\times n}$ is a square, invertible matrix and $X\sim\mathcal{N}(\bm{0},P\bm{I}_{n})$.
 This setting preserves the Gaussian structure while introducing nontrivial correlation through the linear mixing due to $A$.
 Proposition 3 (Closed forms for linear Gaussian channels).
-Let $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$, $A\in\mathbb{R}^{n\times n}$ nonsingular, and $Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n})$, independent. Then for all $t>0$:
-$I(X;Y_{t})$ $=\tfrac{1}{2}\log\det\!\Bigl(I_{n}+\tfrac{P}{t}AA^{\top}\Bigr),$ (35)
-$\mathrm{mmse}(t)$ $=\mathrm{tr}\!\left(\Bigl(P^{-1}I_{n}+\tfrac{1}{t}A^{\top}A\Bigr)^{-1}\right),$ (36)
-$J(Y_{t})$ $=\mathrm{tr}\!\Bigl((PAA^{\top}+tI_{n})^{-1}\Bigr).$ (37)
-Sketch of proof:.
-Since $AX$ is Gaussian with covariance $PAA^{\top}$, we have $Y_{t}\sim\mathcal{N}(\mathbf{0},PAA^{\top}+tI_{n})$.
+Let $X\sim\mathcal{N}(\bm{0},P\bm{I}_{n})$, $\bm{A}\in\mathbb{R}^{n\times n}$ nonsingular, and $Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n})$, independent. Then for all $t>0$:
+$I(X;Y_{t})$ $=\tfrac{1}{2}\log\det\!\Bigl(I_{n}+\tfrac{P}{t}\bm{A}\bm{A}^{\top}\Bigr),$ (36)
+$\mathrm{mmse}(t)$ $=\mathrm{tr}\!\left(\Bigl(P^{-1}I_{n}+\tfrac{1}{t}\bm{A}^{\top}\bm{A}\Bigr)^{-1}\right),$ (37)
+$J(Y_{t})$ $=\mathrm{tr}\!\Bigl((P\bm{A}\bm{A}^{\top}+t\bm{I}_{n})^{-1}\Bigr).$ (38)
+Proof:
+Since $AX$ is Gaussian with covariance $P\bm{A}\bm{A}^{\top}$, we have $Y_{t}\sim\mathcal{N}(\bm{0},P\bm{A}\bm{A}^{\top}+t\bm{I}_{n})$.
 The mutual information (36) follows from the determinant formula for Gaussian differential entropy [2, 14].
-The Fisher information (37) is obtained directly from the Gaussian score function $s_{Y_{t}}(\bm{y})=-(PAA^{\top}+tI_{n})^{-1}\bm{y}$.
+The Fisher information (38) is obtained directly from the Gaussian score function $s_{Y_{t}}(\bm{y})=-(P\bm{A}\bm{A}^{\top}+t\bm{I}_{n})^{-1}\bm{y}$.
 The MMSE formula (37) comes from the linear MMSE estimator of $X$ given $Y_{t}$.
@@
 evaluated numerically.
-In the above algorithm, we consider two alternatives for the DSM learning:
+It should be noted that if the probability density $p_{Y_{t}}$ of the channel output $Y_{t}$ is known, and its score function $s_{Y_{t}}(y)=\nabla_{y}\log p_{Y_{t}}(y)$ is available in closed form (e.g., in the case of an identity function $f$ with a Gaussian Mixture Model (GMM) input $X$), then Step 3 (DSM learning) becomes unnecessary.
+In such cases, one can directly use this analytical score function in Step 4 to perform the Monte Carlo estimation of the Fisher information.
+This approach eliminates the score approximation error while retaining the advantages over conventional methods in terms of computational cost and high-dimensional scalability.
+However, the primary focus of this paper is on more general nonlinear channels where such an analytical solution is intractable or unavailable.
+in Algorithm 1, we consider two alternatives for the DSM learning:
   • Scheme A (per-$t$ training). For each fixed $t$, minimize
@@
 Therefore, the same tail correction can be applied as a universal approximation, with the understanding that higher-order terms may depend on the non-Gaussianity of $X$.
+IV-C Stein Calibration
+In practice, the learned score function $s_{\theta}(\mathbf{y})$ may exhibit a scaling bias.
+To mitigate this, we apply Stein calibration to the score estimates.
+Leveraging Stein’s identity, we compute a scalar correction factor $c$ such that the calibrated score $\tilde{s}_{\theta}(\mathbf{y})=cs_{\theta}(\mathbf{y})$ satisfies the empirical moment constraint $\mathbb{E}[\mathbf{Y}_{t}^{\top}\tilde{s}_{\theta}(\mathbf{Y}_{t})]=-n$, assuming centered data.
+The estimated Fisher information is then computed using this calibrated score.
+IV-D Alternative Score Estimation via Velocity Fields
+While we employed DSM to estimate the score function, the SFB framework is inherently estimator-agnostic.
+Recent generative paradigms, such as flow matching [24] and rectified flow [25], learn a velocity field that drives the probability flow, typically along straight-line trajectories.
+Since the score function can be mathematically derived from the learned velocity field (given the specific scaling factor), these methods offer a promising alternative route for score estimation.
+Leveraging such techniques could potentially improve the numerical stability of MI estimation, particularly in regimes where DSM struggles with curvature in the diffusion path.
+See Appendix G for more details.
 V Nonlinear channel
@@
 $I(X;Y_{t})=I(W;Y_{t}).$ (47)
-Proof.
+Proof:
 By the chain rule of mutual information, we have
@@
 V-B Score-based estimation method for nonlinear channels
-In the following discussin, we assume that the evalution of $f$ is computationally feasible.
+In the following discussion, we assume that the evaluation of $f$ is computationally feasible.
 Namely, sampling from $p_{W}$ is feasible.
@@
 The estimation procedure is still a single-loop procedure, which does not require the access to the posterior density $p_{X|Y_{t}}(\cdot|\cdot)$.
+V-C Error Analysis and Convergence Rates
+This section provides a qualitative analysis of the estimation error for the proposed estimator $\hat{I}$. The total estimation error can be decomposed into four distinct components:
+$|\hat{I}-I|\;\leq\;\varepsilon_{\mathrm{score}}+\varepsilon_{\mathrm{MC}}+\varepsilon_{\mathrm{quad}}+\varepsilon_{\mathrm{tail}}.$
+Each component behaves as follows under standard regularity assumptions:
+  1. 1.
+    Score Approximation Error ($\varepsilon_{\mathrm{score}}$): This term arises from the imperfect learning of the score function $s_{\theta}\approx\nabla\log p_{t}$.
+It scales with the square root of the Fisher divergence loss.
+Improving the neural network capacity and training budget reduces this bias.
+  2. 2.
+    Monte Carlo Variance ($\varepsilon_{\mathrm{MC}}$): Since the Fisher information $J(Y_{t})$ is estimated via sample averaging, this stochastic error follows the standard Central Limit Theorem scaling of $O(N^{-1/2})$, where $N$ is the batch size per grid point.
+  3. 3.
+    Quadrature Error ($\varepsilon_{\mathrm{quad}}$): The discretization of the integral over $u=\log t$ using the trapezoidal rule introduces a deterministic error.
+On a uniform grid with $M$ points (step size $\Delta u\propto M^{-1}$), this error decays rapidly as $O(M^{-2})$.
+  4. 4.
+    Truncation Error ($\varepsilon_{\mathrm{tail}}$): Restricting the integration to $t_{\min},t_{\max}]$ introduces bias from the missing tails.
+The error typically decays as $O(t_{\max}^{-1})$ for the large-$t$ tail and polynomially (e.g., $O(t_{\min}^{\beta})$) for the small-$t$ tail.
+To minimize the total error efficiently, one must balance these components.
+Since the quadrature error $\varepsilon_{\mathrm{quad}}\sim O(M^{-2})$ decays much faster than the Monte Carlo error $\varepsilon_{\mathrm{MC}}\sim O(N^{-1/2})$, a relatively coarse grid $M$ is often sufficient compared to the sample size $N$.
+Theoretically, balancing the rates suggests a scaling of $M\asymp N^{1/5}$.
+In practice, we recommend fixing a sufficiently wide integration range $t_{\min},t_{\max}]$ to suppress $\varepsilon_{\mathrm{tail}}$, and then prioritizing a large batch size $N$ to mitigate the dominant Monte Carlo variance.
 VI Numerical experiments
@@
 To validate the proposed score-based estimation method, we first consider the Gaussian input case where exact closed-form expressions are available.
-Specifically, we consider the Gaussian input case $X\sim\mathcal{N}(0,PI_{n})$ with the AWGN channel model $Y_{t}=X+\sqrt{t}Z$, where $Z\sim\mathcal{N}(0,I_{n})$.
+Specifically, we consider the Gaussian input case $X\sim\mathcal{N}(\bm{0},P\bm{I}_{n})$ with the AWGN channel model $Y_{t}=X+\sqrt{t}Z$, where $Z\sim\mathcal{N}(\bm{0},\bm{I}_{n})$.
 VI-A1 Experimental Setup
@@
 To validate the method beyond the standard AWGN case, we consider the linear Gaussian channel:
-$Y_{t}={\bm{A}}X+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (48)
-where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$ with $P=1.0$, $n=4$.
+$Y_{t}={\bm{A}}X+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n}),$ (49)
+where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $X\sim\mathcal{N}(\bm{0},P\bm{I}_{n})$ with $P=1.0$, $n=4$.
 We train separate score networks (3 hidden layers, 128 units each) for $M=10$ logarithmically spaced noise levels in $t\in[P/200,50P]$.
@@
 To demonstrate the effectiveness of the proposed method for nonlinear channels where closed-form solutions are unavailable, we consider a composite channel with linear mixing followed by elementwise tanh nonlinearity:
-$Y_{t}=\tanh({\bm{A}}X)+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (49)
+$Y_{t}=\tanh({\bm{A}}X)+Z_{t},\quad Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n}),$ (50)
 where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $\tanh(\cdot)$ is applied elementwise.
@@
 Acknowledgments
-This work was supported by JST, CRONOS, Japan Grant Number JPMJCS25N5.
+This work was supported by JST, CRONOS, Japan Grant Number JPMJCS25N5. The author would like to thank Satoshi Takabe for helpful discussions and comments on the first draft of this paper.
+Appendix
+This appendix provides supplementary derivations and score-based integral identities that support (but are not required to follow) the main development of the Score-to-Fisher bridge.
+In particular, we collect concise score-centric representations of entropy, KL divergence, and (normalized) likelihoods, and include optional connections to flow-based generative modeling and optimal transport as additional perspectives.
+VII-A Derivation of Entropy via Thermodynamic Integration
+The relationship between differential entropy and Fisher information is established by the de Bruijn’s identity.
+Let $X_{t}=X+\sqrt{t}Z$ be the diffusion process where $Z\sim\mathcal{N}(\bm{0},\bm{I}_{n})$.
+The time derivative of the entropy $h(X_{t})$ is given by:
+$\frac{d}{dt}h(X_{t})=\frac{1}{2}J(X_{t}),$ (52)
+where $J(X_{t})$ is the Fisher information.
+To evaluate the entropy $h(X)$ of the clean data (at $t=0$), we compare the data process $X_{t}$ with a reference Gaussian process $X_{G,t}$ starting from a standard Gaussian distribution.
+Integrating the difference between their derivatives from $t=0$ to $T$ yields:
+$h(X_{T})-h(X_{G,T})$ $=(h(X)-h(X_{G,0}))$
+$+\frac{1}{2}\int_{0}^{T}(J(X_{t})-J(X_{G,t}))dt.$ (53)
+As $T\to\infty$, both $X_{T}$ and $X_{G,T}$ are dominated by the added Gaussian noise and the entropy difference, $h(X_{T})-h(X_{G,T})$, vanishes (under mild moment/regularity conditions).
+Rearranging the terms and taking the limit $T\to\infty$, we obtain the integral representation:
+$h(X)=h(X_{G,0})-\frac{1}{2}\int_{0}^{\infty}\left(J(X_{t})-J(X_{G,t})\right)dt.$ (54)
+Assuming a standard Gaussian reference $X_{G,0}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$, we have $h(X_{G,0})=\frac{n}{2}\log(2\pi e)$.
+Furthermore, since $X_{G,t}\sim\mathcal{N}(\bm{0},(1+t)\bm{I}_{n})$, the Fisher information of the reference process is given exactly by $J(X_{G,t})=\frac{n}{1+t}$.
+Substituting this into the integral yields the exact representation:
+$h(X)=\frac{n}{2}\log(2\pi e)-\frac{1}{2}\int_{0}^{\infty}\left(J(X_{t})-\frac{n}{1+t}\right)dt.$ (55)
+VII-B Derivation of KL Divergence Representation
+Let $p_{t}$ and $q_{t}$ denote the probability densities of the random variable ${X}_{t}={X}+\sqrt{t}{Z}$ where ${Z}\sim\mathcal{N}(\bm{0},\bm{I})$, derived from initial distributions $p$ and $q$, respectively.
+Both $p_{t}$ and $q_{t}$ satisfy the heat equation $\partial_{t}\rho=\frac{1}{2}\Delta\rho$.
+Consider the time derivative of the KL divergence $D_{\text{KL}}(p_{t}\|q_{t})=\int p_{t}\log(p_{t}/q_{t})d\bm{x}$:
+$\frac{d}{dt}D_{\text{KL}}(p_{t}\|q_{t})$ $=\int\frac{\partial p_{t}}{\partial t}\log\frac{p_{t}}{q_{t}}d\bm{x}+\int p_{t}\left(\frac{\partial_{t}p_{t}}{p_{t}}-\frac{\partial_{t}q_{t}}{q_{t}}\right)d\bm{x}$
+$=\frac{1}{2}\int\Delta p_{t}\log\frac{p_{t}}{q_{t}}d\bm{x}+\underbrace{\frac{1}{2}\int\Delta p_{t}d\bm{x}}_{0}$
+$\quad-\frac{1}{2}\int\frac{p_{t}}{q_{t}}\Delta q_{t}d\bm{x}.$ (56)
+Using integration by parts (assuming boundary terms vanish), we obtain:
+$\frac{d}{dt}D_{\text{KL}}(p_{t}\|q_{t})$
+$=-\frac{1}{2}\int\nabla p_{t}\cdot\nabla\log\frac{p_{t}}{q_{t}}d\bm{x}+\frac{1}{2}\int\nabla\left(\frac{p_{t}}{q_{t}}\right)\cdot\nabla q_{t}d\bm{x}$
+$=-\frac{1}{2}\int p_{t}(\bm{s}_{p}-\bm{s}_{q})\cdot(\bm{s}_{p}-\bm{s}_{q})d\bm{x}$
+$=-\frac{1}{2}\mathbb{E}_{p_{t}}[\|\bm{s}_{p}(\bm{x},t)-\bm{s}_{q}(\bm{x},t)\|^{2}].$ (57)
+Integrating from $t=0$ to $\infty$, and assuming $D_{\text{KL}}(p_{\infty}\|q_{\infty})=0$ (since both distributions converge to the same noise distribution), we arrive at the representation:
+$D_{\text{KL}}(p\|q)=\frac{1}{2}\int_{0}^{\infty}\mathbb{E}_{p_{t}}\left[\|\bm{s}_{p}(\bm{x},t)-\bm{s}_{q}(\bm{x},t)\|^{2}\right]dt.$ (58)
+This identity implies that the KL divergence is equivalent to the time integral of the Fisher divergence along the diffusion path.
+Intuitively, the global information difference between $p$ and $q$ is obtained by accumulating the local geometric discrepancies of their score functions throughout the smoothing process.
+VII-C Exact Log-Likelihood Ratio via Path Integration
+Since the score function is the gradient of the log-density, $\bm{s}(\bm{x})=\nabla_{\bm{x}}\log p_{X}(\bm{x})$, the exact log-likelihood ratio between any target state $\bm{x}$ and a reference state $\bm{x}_{0}$ can be computed by line integration along a path $\gamma$:
+$\log\frac{p_{X}(\bm{x})}{p_{X}(\bm{x}_{0})}=\int_{\bm{x}_{0}}^{\bm{x}}\bm{s}(\bm{u})\cdot d\bm{u}.$ (59)
+This formulation allows for the evaluation of relative probabilities (or energy differences) without knowledge of the intractable partition function $Z$.
+This property is particularly powerful for applications such as hypothesis testing, anomaly detection, and MCMC transition acceptance, where only probability ratios are required.
+For high-dimensional data, radial integration from the dataset centroid is a practical choice to ensure numerical stability.
+VII-D Recovery of the Partition Function
+Consider an energy-based representation $p_{X}(\bm{x})=e^{-\phi(\bm{x})}/Z$, where $\phi(\bm{x})$ is the recovered potential (energy) and $Z$ is the unknown partition function.
+By taking the expectation of the log-density $\log p_{X}(\bm{x})=-\phi(\bm{x})-\log Z$ with respect to the data distribution, we have
+$h(X)=-\mathbb{E}_{X}[\phi(X)]-\log Z.$ (60)
+Solving for $\log Z$ yields the log partition function:
+$\log Z=h(X)-\mathbb{E}_{X}[\phi(X)].$ (61)
+Here, $\mathbb{E}[\phi(X)]$ is computed via sample averaging over the dataset.
+From a thermodynamic perspective, this relationship corresponds to the definition of the Helmholtz free energy, $\mathcal{F}=\mathcal{U}-\mathcal{S}$ (assuming unit temperature).
+By identifying the internal energy as $\mathcal{U}=\mathbb{E}[\phi(X)]$ and the entropy as $\mathcal{S}=h(X)$, the log-partition function is equivalent to the negative free energy, $\log Z=\mathcal{S}-\mathcal{U}=$-$\mathcal{F}$.
+The proposed Score-to-Fisher bridge opens a promising avenue for scalable partition function estimation in energy-based models (EBMs).
+Traditionally, evaluating the partition function ZZ—or equivalently the free energy—has been a major computational bottleneck, often requiring expensive MCMC simulations like annealed importance sampling.
+However, our framework allows for the direct computation of the log-partition function via the identity $\log Z=h(X)-\mathbb{E}[\phi(X)]$, where the entropy $h(X)$ is efficiently estimated by the Fisher integral.
+This approach effectively converts the integration problem over a high-dimensional volume into a tractable one-dimensional integration along a diffusion path.
+VII-E Exact Log-Likelihood
+Substituting (61) back into the density definition, the exact log-likelihood for any query sample $\bm{x}$ (including test data) is given by
+$\log p_{X}(\bm{x})=-\phi(\bm{x})+\mathbb{E}_{X}[\phi(X)]-h(X).$ (62)
+This formulation allows for exact likelihood evaluation and model comparison without explicit integration over the entire state space, provided the score function is accurately learned.
+VII-F Unified Representation of Information-Theoretic Quantities
+In this work, we have primarily focused on the estimation of mutual information via the Score-to-Fisher bridge.
+However, the proposed framework—grounded in the interplay between the local score function $\nabla_{\bm{x}}\log p(\bm{x})$ and the Fisher information—offers a unified perspective for various fundamental quantities in information theory.
+Table I summarizes the correspondence between classical Shannon definitions and their score-based integral representations derived in this appendix.
+A key insight here is the transformation of computational complexity: intractable global volume integrations (e.g., for entropy and partition functions) are converted into tractable estimations of local geometry (scores) followed by 1D integrations.
+This “Score-Centric” view provides a consistent and scalable foundation for high-dimensional statistical modeling beyond mutual information.
+TABLE I: Information-Theoretic Quantities via Score-Based Representations. This framework translates global integration problems into local score estimation and 1D integration.
+Quantity Definition Score-Based Representation
+Differential Entropy $h(X)=-\mathbb{E}[\log p]$ $\frac{n}{2}\log(2\pi e)-\frac{1}{2}\int_{0}^{\infty}\left(J(X_{t})-\frac{n}{1+t}\right)dt$
+Mutual Information $I(X;Y_{T})=h(Y_{T})-h(Y_{T}|X)$ $\frac{1}{2}\int_{T}^{\infty}\left(\frac{n}{t}-J(Y_{t})\right)dt$
+KL Divergence $D_{\text{KL}}(p\|q)=\mathbb{E}_{p}[\log\frac{p}{q}]$ $\frac{1}{2}\int_{0}^{\infty}\mathbb{E}_{p_{t}}\left[\|{s}_{p}(\bm{x},t)-{s}_{q}(\bm{x},t)\|^{2}\right]dt$
+Log-Likelihood Ratio $\log\frac{p(\bm{x})}{p(\bm{x_{0}})}$ $\int_{\bm{x}_{0}}^{\bm{x}}{s}_{p}(\bm{u})\cdot d\bm{u}$  (Line Integral)
+Log-Partition Function $\log Z=\log\int e^{-\phi(\bm{x})}d\bm{x}$ $h(X)-\mathbb{E}_{X}\left[\phi(X)\right]$  (Free Energy)
+Likelihood $\log p(\bm{x})$ $\phi(\bm{x})-\displaystyle h(X)+\mathbb{E}_{X}\left[\phi(X)\right]$
+VII-G Score Estimation from Velocity Fields in Rectified Flow
+Recent generative models often employ rectified flow [25] (or flow matching [24]), which defines a probability path via linear interpolation between data $\bm{x}\sim p_{X}$ and noise $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$:
+$\bm{x}_{\tau}=(1-\tau)\bm{x}+\tau\bm{z},\quad\tau\in[0,1].$ (63)
+Instead of learning the score function directly, these models learn a velocity field $\bm{v}_{\tau}(\bm{x}_{\tau})$ that minimizes the regression loss against the target velocity $\bm{u}_{\tau}=\bm{z}-\bm{x}$.
+The optimal velocity field is given by the conditional expectation $\bm{v}_{\tau}(\bm{x}_{\tau})=\mathbb{E}[\bm{z}-\bm{x}|\bm{x}_{\tau}]$.
+To apply the Score-to-Fisher bridge, we need to recover the score function $\nabla_{\bm{x}_{\tau}}\log p_{\tau}(\bm{x}_{\tau})$ from the learned velocity $\bm{v}_{\tau}$.
+Leveraging Tweedie’s formula adapted for the linear interpolation path, the posterior expectations of the signal and noise are related to the score function by:
+$\hat{\bm{x}}=\frac{\bm{x}_{\tau}+\tau^{2}\nabla\log p_{\tau}}{1-\tau},\quad\hat{\bm{z}}=-\tau\nabla\log p_{\tau}.$ (64)
+Substituting these into the relation $\bm{v}_{\tau}=\hat{\bm{z}}-\hat{\bm{x}}$, we obtain the explicit connection between the velocity and the score:
+$\bm{v}_{\tau}(\bm{x}_{\tau})=-\frac{\bm{x}_{\tau}+\tau\nabla_{\bm{x}_{\tau}}\log p_{\tau}(\bm{x}_{\tau})}{1-\tau}.$ (65)
+Inverting this equation yields the score function in terms of the velocity field:
+$\nabla_{\bm{x}_{\tau}}\log p_{\tau}(\bm{x}_{\tau})=-\frac{\bm{x}_{\tau}+(1-\tau)\bm{v}_{\tau}(\bm{x}_{\tau})}{\tau}.$ (66)
+Furthermore, to map this to the standard additive noise model ${Y}_{\sigma}=\bm{x}+\sigma\bm{z}$ used in our Fisher integral (where $\sigma=\tau/(1-\tau)$), the score scales as $\nabla_{{Y}}\log p_{Y}({Y})=(1-\tau)\nabla_{{X}}\log p_{\tau}({X})$.
+This derivation confirms that flow-based models can serve as plug-in estimators for the Fisher information and, consequently, the mutual information.
+VII-H Optimal Transport Cost via Score Integration
+The same score-to-velocity identity also provides an optimal transport(OT)-flavored diagnostic.
+In particular, plugging the induced velocity into the Benamou-Brenier functional yields an upper bound on the Wasserstein-2 distance between the data and the Gaussian prior, offering a geometric measure of how far the data are from the Gaussian prior.
+The proposed framework relates to optimal transport theory through the dynamic formulation of the Wasserstein-2 distance.
+According to the Benamou–Brenier formula, the squared Wasserstein distance $W_{2}^{2}(p,q)$ between two distributions equals the infimum of the kinetic energy $\int_{0}^{1}\mathbb{E}_{\rho_{\tau}}\!\left[\|\bm{v}_{\tau}(\bm{x})\|^{2}\right]d\tau$ over all admissible probability paths $\{\rho_{\tau}\}_{\tau\in[0,1]}$ and velocity fields $\{\bm{v}_{\tau}\}$ satisfying the continuity equation $\partial_{\tau}\rho_{\tau}+\nabla\!\cdot(\rho_{\tau}\bm{v}_{\tau})=0$.
+In the context of rectified flow, we consider the linear interpolation path ${X}_{\tau}=(1-\tau){X}+\tau{Z}$ connecting the data distribution $p_{X}$ and the Gaussian prior $p_{Z}=\mathcal{N}(\bm{0},\bm{I})$.
+Let $p_{\tau}$ denote the marginal density of $\bm{X}_{\tau}$.
+Using the identity derived from this linear structure, the corresponding velocity field can be expressed in terms of the score as
+$\bm{v}_{\tau}(\bm{x})\;=\;-\frac{\bm{x}+\tau\,\nabla_{\bm{x}}\log p_{\tau}(\bm{x})}{1-\tau},\qquad\bm{x}\in\mathbb{R}^{n},\ \tau\in(0,1).$ (67)
+Therefore, for this specific admissible path $(p_{\tau},\bm{v}_{\tau})$, the Benamou–Brenier functional yields an upper bound on the Wasserstein cost:
+$W_{2}^{2}(p_{X},p_{Z})\;\leq\;\int_{0}^{1}\mathbb{E}_{p_{\tau}}\left[\left\|\frac{\bm{x}_{\tau}+\tau\nabla_{\bm{x}_{\tau}}\log p_{\tau}(\bm{x}_{\tau})}{1-\tau}\right\|^{2}\right]d\tau.$ (68)
+Equality holds if and only if the chosen path coincides with the optimal displacement interpolation (i.e., the $W_{2}$-geodesic) between $p_{X}$ and $p_{Z}$.
+Rectified flow models aim to learn a velocity field consistent with such an optimal transport path, in which case the above upper bound becomes tight.